### v1

# Learning Contracts in Hierarchical Multi-Agent Systems

Abstract
The emergence of Machine Learning systems everywhere raises new challenges, such as dealing with interactions or competition between multiple learners. In that goal, we study multi-agent sequential decision-making by considering principal-agent interactions in a tree structure. In this problem, the reward of a player is influenced by the actions of her children, who are all self-interested and non-cooperative. Our main finding is that it is possible to steer all the players towards the globally optimal set of actions by simply allowing single-step contracts between them. A contract is established between a principal and one of her agents: the principal actually offers the proposed payment if the agent picks the recommended action. The analysis poses specific challenges due to the intricate interactions between the nodes of the tree. Within a bandit setup, we propose algorithmic solutions for the players to end up being no-regret with respect to the optimal pair of actions and contracts. In the long run, allowing contracts makes the players act as if they were collaborating together, although they remain non-cooperative.
Machine Learning, ICML
1 Introduction
The increasing deployment of Machine Learning (ML) systems in real-world environments introduces specific challenges, particularly regarding interactions and competition among self-interested learning agents. This has led to a growing literature that incorporates economic considerations into ML settings. A large part of these works focuses on mechanism design (Laffont & Martimort, 2009; Nisan & Ronen, 1999; Myerson, 1989), where a principal (she) wants an agent (he) to perform a specific task and needs to incentivize the agent so he participates. Usually, the mechanism is a payment based on the agent’s behavior. The majority of this literature assumes that one principal interacts with one agent or more. Unfortunately, real-world scenarios generally present much more intricate structures. When an agent accepts a contract from a principal, it is likely that this one agent then delegates a part of his effort to another agent, and so on and so forth. Extending this reasoning reveals a natural hierarchy of players, forming a chain of interactions through principal-agent relations. It raises an issue about the behavior of players organized in a general structure of relations, who interact together through transfers in a repeated game; put differently:
Can self-interested learning agents converge to social welfare maximization through online contracting with their neighbors?
Although the analysis of equilibrium in learning games has been extensively studied (Wang & Sandholm, 2002; Celli et al., 2020; Chakrabarti et al., 2024; Daskalakis et al., 2024), these settings solely involve one principal, her agents, and sometimes a central planner. As a first step towards solving our question in a general decentralized structure with no mediator, we assume an intricate principal-agent structure of the players, organized as a tree (such structures are often studied in games, see, Zhang & Sandholm, 2022; Fiegel et al., 2023). The players sequentially take actions on a bandit instance, with their action influencing both their own reward and the reward of their principal. Our objective is to analyze the convergence of the entire set of players towards the global optimum, defined as the set of actions that maximizes the sum of all utilities. However, such convergence cannot occur endogenously, as individual players have no incentive to account for their principal’s reward in their decision-making. To resolve this, we introduce contracts (learned over time) from the principal to the agent, that help to align the players’ utilities. Contract Theory often presents two primary challenges: Hidden Action (a.k.a. Moral Hazard, see, Mirrlees, 1999), which refers to the scenario where the principal cannot observe the actions taken by the agent; and Hidden Type, also known as Adverse Selection, which refers to the scenario where the principal cannot observe the agent’s type. We specifically focus on the latter in this work.
Consider the illustrative example of Federated Learning (Kairouz et al., 2021; Li et al., 2020), which was introduced due to the impossibility of processing tasks with huge amounts of data on one central server - due to privacy constraints, network bandwidth, etc… It motivates the emergence of FL, where the collection and the processing of data is decentralized on the nodes of a network. However, in practice, agents must be incentivized to participate, which introduces the need for contract theory (see, Zhan et al., 2020, 2021; Karimireddy et al., 2022; Liu et al., 2022; Capitaine et al., 2024, for contract theory and FL). Now, consider a FL setting where a principal pays agents to perform a learning task, which we model as selecting an arm in a bandit instance. These agents, in turn, may delegate part of their task to data servers, offering them contracts to collect and process data—which again corresponds to choosing an action in a bandit instance. This cascading delegation naturally leads to a hierarchical structure of decision-making and incentives. This example highlights the need to develop models that handle complex, multi-layered contracting mechanisms in Federated Learning. While prior works studied the delegation of learning tasks between a principal and an agent through the lens of contract theory (see, e.g., Ananthakrishnan et al., 2024a; Saig et al., 2024b), they do not go beyond the case of one principal and agents.
Thus, we present a general formulation of a game between players organized as a tree, that is versatile enough to encompass a variety of applications, including the example of Federated Learning that we just mentioned. At each step, a node (a player) observes the contract (an action and an associated payment) offered by her parent. Then, the player takes an action on her multi-armed bandit instance and offers a contract to her children. The reward of a node depends on her action as well as the actions taken by her children.
Contributions. After introducing this original setting, we introduce COBRA: Contracting Online with Bandit Rewards and several Agents, an algorithm that enables players to learn optimal contracts before executing a bandit subroutine on the updated problem instance, which accounts for the estimated contracts. Each player must wait until their agents have sufficiently learned and are best-responding with enough accuracy before initiating the process of learning contracts. Finally, we derive regret bounds for the players as well as an upper bound for the social welfare regret that evolves as $o(T)$, when every player follows COBRA.
From a technical perspective, the regret bounds result are achieved by decomposing the regret of each player into three components: the error due to her own suboptimal actions, the error arising from inaccuracies in the contract payments, and the error caused by the suboptimal decisions taken by her agents. Interestingly, by controlling individual regrets while accounting for contracts, we are able to manage the collective regret of all players, effectively treating them as if they were collaborating. It reveals that it is possible to orchestrate selfish players without the need for a central server, solely by enabling contracts.
A large part of the regret is due to suboptimal actions picked by the children of a node, which leads to a propagation of the regret in the tree, which grows as we approach from the root. However, the regret propagation is controlled through the decomposition and the formulation of 4.3 for the action regret only, so it does not explode.
2 Related Works
The study of social welfare (also referred to as global utility) in interactions between a principal and an agent (Grossman & Hart, 1992; Bendor et al., 2001) has been a long-standing area of research in the mechanism design literature that recently benefited from data-driven approaches (Bergemann et al., 2024) and, more recently, applications of large language models (Duetting et al., 2024; Dubey et al., 2024). In this work, we focus specifically on the problem of learning optimal contracts in an online setting (see, e.g., Zuo, 2024a, b). The study of learning through sequential interactions has also gained attention (Han et al., 2024; Kolumbus et al., 2024), especially with the recent rise of algorithmic contract design (Dütting et al., 2022; Bacchiocchi et al., 2023; Guruganesh et al., 2024; Bollini et al., 2024; Collina et al., 2024a). For an accessible introduction to algorithmic contract design, see Dütting et al. (2024). While prior work has examined interactions between a principal and multiple agents (Dütting et al., 2023), to the best of our knowledge, this paper is the first to address the more complex scenario of a nested chain of interacting principals and agents. From a statistical point of view, our work draws heavily on ideas from the bandit literature (Lattimore & Szepesvári, 2020; Slivkins et al., 2019), though it diverges significantly from the literature on multi-player or competing bandits (Boursier & Perchet, 2022; Liu et al., 2020), which typically considers simultaneous actions without transfers.
The emergence of large models considered as economic agents (Shapira et al., 2024; Immorlica et al., 2024; Deng et al., 2024) motivates the study of learning in repeated Stackelberg games (Brown et al., 2024; Morri et al., 2025). The latter is challenging because the leader generally does not know the follower’s preferences (Arunachaleswaran et al., 2024). From a global perspective, welfare maximization with learning agents at stake is a complex task (Anagnostides et al., 2024): a central question is whether players can converge to equilibrium, which has been explored in Zhang et al. (2023). However, their setup assumes the presence of a mediator interacting with multiple agents, thereby overlooking the intricate interactions and decentralized nature of the problem. Incentives have proved to be efficient to align players with initially misaligned goals (Saig et al., 2024a; Maheshwari et al., 2024; Freund & Hssaine, 2025).
Finally, our work heavily relies on the interactions between principal-agent bandit learners, as they are introduced in Scheid et al. (2024a, b), but their approach differs in that the players do not exhibit an intricate hierarchical structure (which raise most of the technical challenges here). The concept of interacting bandit players has been first introduced in Dogan et al. (2023b, a), with notable extensions to Markov Decision Processes (Ivanov et al., 2024; Ben-Porat et al., 2024; Wu et al., 2024) and frameworks where agents are constrained in probability to stay nearby their best responses (Liu & Ratliff, 2024).
3 Setting
Figure 1: Illustration of the game structure: each node of the tree (in red) represents a player and the blue edges correspond to the principal/agent interactions between the players.
Model. We consider a set of of players $\mathrm{V}$ organized as a tree $\mathcal{T}=(\mathrm{V},\mathrm{E})$ where $\mathrm{E}$ represents the set of edges between the players. These players engage in an online learning game over a horizon of $T$ steps. At each time step $t\in\{1,\ldots,T\}$, each player $\mathrm{v}\in\mathrm{V}$ selects an action $A_{t}^{\mathrm{v}}$ from a common action set $\mathcal{A}$ of size $K$ (it is equivalent to consider the same action set or distinct action sets for the players).
We assume that players are expected-utility maximizers and do not strategically manipulate other players in the long run. Although it is possible for a learning agent to achieve its Stackelberg value through strategic interactions (Goktas et al., 2022; Zhao et al., 2023; Collina et al., 2024b; Haghtalab et al., 2024), this is not always the case (as shown in Ananthakrishnan et al., 2024b). Furthermore, we emphasize that our focus is not on achieving a Stackelberg optimum in this work.
A key distinction from existing literature is that the players in our framework are organized in a hierarchical tree structure with depth $\mathrm{D}$ and breadth $\mathrm{B}$, reflecting principal-agent relationships. Each player $\mathrm{v}$ (represented as a node in the tree) serves as the principal for her children, denoted $\mathrm{C}(\mathrm{v})$, while simultaneously acting as an agent for her parent, $\mathrm{P}(\mathrm{v})$. Without loss of generality, we assume that each player has exactly $\mathrm{B}$ children. The tree is indexed by its depth $d\in\{1,\ldots,\mathrm{D}\}$, where $d=1$ corresponds to the leaf nodes, and $d=\mathrm{D}$ represents the root. The total number of vertices in the tree is $\text{Card}(\mathrm{V})=\sum_{k=1}^{\mathrm{D}}\mathrm{B}^{k}=(\mathrm{B}^{\mathrm{D}+1}-1)/(\mathrm{B}-1)$. For clarity, we refer to the specific player under consideration as $\mathrm{v}$, her children as $\mathrm{w}$, and her parent as $\mathrm{u}$ throughout the analysis.
Reward. Within this tree structure, the reward of a player $\mathrm{v}$ at time $t$ depends on her action $A_{t}^{\mathrm{v}}$ and the actions of all her children, denoted as $A_{t}^{\mathrm{C}(\mathrm{v})}=(A_{t}^{\mathrm{w}})_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}$. The reward of player $\mathrm{v}$ at round $t$ is given by $X_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})$, which is drawn at random following:
$X_{t}^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})=\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})+z_{t}^{\mathrm{v}}\;,$ (1)
where $(z_{t}^{\mathrm{v}})_{t\in\mathbb{N}^{\star}}$ are i.i.d. zero-mean sub-Gaussian random variables, and $(\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})}))_{(a,a^{\mathrm{C}(\mathrm{v})})\in\mathcal{A}^{\mathrm{B}+1}}\in\mathbb{R}^{K^{\mathrm{B}+1}}$. For any $(a,a^{\mathrm{C}(\mathrm{v})})\in\mathcal{A}^{\mathrm{B}+1}$, the expected reward is given by:
$\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})=\mathbb{E}[X_{t}^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})],$
which represents the mean reward of player $\mathrm{v}$ when the actions $(a,a^{\mathrm{C}(\mathrm{v})})$ are played. The quantities $(\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})}))_{(a,a^{\mathrm{C}(\mathrm{v})})\in\mathcal{A}^{\mathrm{B}+1}}$ are unknown to the players, whose objective is to learn the actions with the highest mean rewards. This is achieved through a sequential selection of actions. We impose the following assumption on the mean rewards, which is equivalent to assuming bounded rewards, as they can always be rescaled.
Assumption 3.1.
For any player $\mathrm{v}\in\mathrm{V}$ and any actions $a\in\mathcal{A},a^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}$, we have that $\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})\in[0,1]$.
On the Necessity of Contracting. In this setup, as in any game, the primary difficulty lies in the fact that a player’s reward depends on the actions of other players. Without additional mechanisms, the global utility of the participants is unlikely to be maximized: some players (e.g., leaf nodes) may achieve reasonably good utility, while others may gain little to no utility (see Scheid et al., 2024a, Section 2 for a detailed proof in a simple two-players example). To address this issue, we allow players to establish contracts with one another.
We define a short-term contract as a pair $(B_{t}^{\mathrm{w}},\tau_{t}(\mathrm{w}))\in\mathcal{A}\times\mathbb{R}_{+}$, where player $\mathrm{v}$ offers a transfer $\tau_{t}(\mathrm{w})$ to her child $\mathrm{w}$ if the latter plays the recommended action $B_{t}^{\mathrm{w}}$. Player $\mathrm{w}$ can either accept the contract and play $B_{t}^{\mathrm{w}}$ to receive the transfer $\tau_{t}(\mathrm{w})$ or reject the contract and play any action his choice, forfeiting the transfer.
Consider player $\mathrm{v}\in\mathrm{V}$. At each step $t\in\{1,\ldots,T\}$, $\mathrm{v}$ first observes the contract $(B_{t}^{\mathrm{v}},\tau_{t}(\mathrm{v}))$ offered by her parent $\mathrm{P}(\mathrm{v})$. Then, $\mathrm{v}$ offers a contract $\mathcal{C}_{t}^{\mathrm{w}}=(B_{t}^{\mathrm{w}},\tau_{t}(\mathrm{w}))$ to any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, which includes recommended actions $B_{t}^{\mathrm{C}(\mathrm{v})}=(B_{t}^{\mathrm{w}})_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}$ and transfers $(\tau_{t}(\mathrm{w}))_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}$. The resulting utility is:
$\mathcal{U}_{t}^{\mathrm{v}}$ $=X_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\cdot\tau_{t}(\mathrm{v})$
$\quad-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{B_{t}^{\mathrm{w}}}(A_{t}^{\mathrm{w}})\cdot\tau_{t}(\mathrm{w})\;,$ (2)
where the utility accounts for the contract received from $\mathrm{P}(\mathrm{v})$. At the end of each round, in addition to observing the bandit feedback $X_{t}^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})$, player $\mathrm{v}$ also observes the actions of her children $(A_{t}^{\mathrm{w}})_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}$.
Challenges. Note that $\mathrm{v}$ also acts as an agent for her parent $\mathrm{P}(\mathrm{v})$ and receives a contract $(B_{t}^{\mathrm{v}},\tau_{t}(\mathrm{v}))$. As a result, $\mathrm{v}$ faces a three-fold trade-off: balancing the contract received from her parent, the action she chooses, and the contracts she offers to her children.
4 Algorithm and Results
Players’ Policies. Consider a player $\mathrm{v}$, whose history is defined as $\mathcal{H}_{0}^{\mathrm{v}}=\varnothing$, and for any $t>0$:
$\mathcal{H}_{t}^{\mathrm{v}}=\mathcal{H}_{t-1}^{\mathrm{v}}\cup\{U_{t}^{\mathrm{v}},B_{t}^{\mathrm{v}},\tau_{t}(\mathrm{v}),A_{t}^{\mathrm{C}(\mathrm{v})},X_{t}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})\},$
where $((U_{t}^{\mathrm{v}})_{t\in\mathbb{N}^{\star}})_{\mathrm{v}\in\mathrm{V}}$ are families of independent uniform random variables in $[0,1]$ used for policy randomization. A policy for player $\mathrm{v}\in\mathrm{V}$ is defined as a mapping:
$\pi^{\mathrm{v}}$ $\colon(U_{t+1}^{\mathrm{v}},\mathcal{H}_{t}^{\mathrm{v}},B_{t+1}^{\mathrm{v}},\tau_{t+1}(\mathrm{v}))$ (3)
$\quad\mapsto(B_{t+1}^{\mathrm{C}(\mathrm{v})},\tau_{t+1}(\mathrm{C}(\mathrm{v})),A_{t+1}^{\mathrm{v}})\;.$
In this setup, player $\mathrm{v}$ must determine her optimal action $A^{\mathrm{v}}$ while simultaneously ensuring that her children $\mathrm{C}(\mathrm{v})$ perform actions $B^{\mathrm{C}(\mathrm{v})}$ that maximize her mean reward $\theta^{\mathrm{v}}(A^{\mathrm{v}},B^{\mathrm{C}(\mathrm{v})})$. To achieve this, the contract $(B^{\mathrm{C}(\mathrm{v})},\tau(\mathrm{C}(\mathrm{v})))$ must incentivize the children to follow the recommended actions. The transfer $\tau(\mathrm{C}(\mathrm{v}))$ poses a critical trade-off: if set too high, it reduces $\mathrm{v}$’s utility; if set too low, the children may ignore the recommendations. Designing such a policy requires sophisticated algorithmic strategies to balance these competing objectives effectively.
Optimal Incentives. For any players $\mathrm{v}\in\mathrm{V}$ and $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, any action $b\in\mathcal{A}$ that $\mathrm{v}$ wants to be played by $\mathrm{w}$, we define the optimal contract $(b,\tau^{\star}_{b}(\mathrm{w}))$ as the minimal transfer in hindsight that $\mathrm{v}$ needs to offer to $\mathrm{w}$ so that $b$ is in expectation the best action that $\mathrm{w}$ could play.
The definition of $\tau^{\star}$ appears by induction starting with the optimal contract for the leafs and then ascending up to the root. Consider a player $\mathrm{v}$ at depth $2$ who interacts with one of her children $\mathrm{w}$ at depth $1$ (a leaf) and assume that $\mathrm{w}$ is best-responding. Since $\mathrm{w}$ is a leaf and has no agent, $\mathrm{C}(\mathrm{w})=\varnothing$. Hence, if $\mathrm{v}$ offers a contract $\mathcal{C}^{\mathrm{w}}=(B^{\mathrm{w}},\tau(\mathrm{w}))$ to $\mathrm{w}$ and wants to make $B^{\mathrm{w}}$ better than any other action, we need to have for any $a\in\mathcal{A}$
$\theta^{\mathrm{w}}(B^{\mathrm{w}})+\text{1}_{B^{\mathrm{w}}}(B^{\mathrm{w}})\,\tau(\mathrm{w})>\theta^{\mathrm{w}}(a)+\text{1}_{B^{\mathrm{w}}}(a)\,\tau(\mathrm{w})\;.$
Thus, for any $\tau>\tau^{\star}_{B^{\mathrm{w}}}(\mathrm{w})$, and any $a\neq B^{\mathrm{w}}$, we must have
$\theta^{\mathrm{w}}(B^{\mathrm{w}})+\tau>\theta^{\mathrm{w}}(a)\;,$
which gives, since we want $\tau^{\star}_{B^{\mathrm{w}}}(\mathrm{w})$ as small as possible
$\tau^{\star}_{B^{\mathrm{w}}}(\mathrm{w})=\max_{a\in\mathcal{A}}\theta^{\mathrm{w}}(a)-\theta^{\mathrm{w}}(B^{\mathrm{w}})\;.$ (4)
Therefore, for a set of actions $(A_{t}^{\mathrm{v}},B^{\mathrm{C}(\mathrm{v})}_{t})$: if $\mathrm{v}$ knows the optimal contracts $(b,\tau^{\star}_{b}(\mathrm{w}))_{b\in\mathcal{A},\mathrm{w}\in\mathrm{C}(\mathrm{v})}$ and her agents are best-responding, she obtains an expected utility at step $t$
$\mathbb{E}[\mathcal{U}_{t}^{\mathrm{v}}]$ $=\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},B^{\mathrm{C}(\mathrm{v})}_{t})+\text{1}_{B^{\mathrm{v}}_{t}}(A^{\mathrm{v}}_{t})\tau_{t}(\mathrm{v})$ (5)
$\quad-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{B^{\mathrm{w}}_{t}}(\mathrm{w})\;.$
Optimization Problem in Hindsight. The quantities in (4) and (5) are straightforward to compute for players at the leaves of the tree, but the task becomes significantly more complex for nodes in the core of the tree. Consider a player $\mathrm{v}\in\mathrm{V}$ at depth $d\geqslant 3$. We assume fully rational players with perfect knowledge of the game, i.e., all children $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ are aware of the optimal average utility $\mu^{\star,\mathrm{w}}(b)$ they can extract from any action $b\in\mathcal{A}$. This utility is well-defined as long as the optimal contracts are specified, as in (4).
For a single round, player $\mathrm{v}$’s objective is to identify a set of actions $(a^{\star},b^{\star,\mathrm{C}(\mathrm{v})})$ and propose an optimal contract $\mathcal{C}^{\star}(b^{\star,\mathrm{C}(\mathrm{v})})=(b^{\star,\mathrm{C}(\mathrm{v})},\tau^{\star}_{b^{\star,\mathrm{C}(\mathrm{v})}}(\mathrm{C}(\mathrm{v})))$ that solves the following optimization problem:
maximize $\text{maximize }\;\theta^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau(\mathrm{w})$ (6)
[…]
for all $\mathrm{w}\in\mathrm{C}(\mathrm{v})$. This problem requires knowledge of the optimal contract $(b^{\mathrm{w}},\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w}))$ for each child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ to be solved exactly. The transfer $\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})$ is defined as the minimum utility that must be transferred from $\mathrm{v}$ to $\mathrm{w}$ to ensure that $\mathrm{w}$ selects $b^{\mathrm{w}}$ as their optimal action in hindsight
Social Welfare is defined as the total reward accumulated across all players. Equation 6 shows that the introduction of contracts incentivizes players to select actions that not only optimize their own rewards, but also contribute to the overall benefit of others. Actions that lead to a higher collective reward thus become less costly, aligning individual incentives with social welfare.
Lemma 4.1.
For any player $\mathrm{v}\in\mathrm{V}$, the optimal contracts $\mathcal{C}^{\star}(b^{\mathrm{C}(\mathrm{v})})=(b^{\mathrm{w}},\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w}))_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}$ for any $b^{\mathrm{C}(\mathrm{v})}=(b^{\mathrm{w}})_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}$ are given by
$\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})=\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})\;,$
and the expected utility for the couple of actions $(a,b^{\mathrm{C}(\mathrm{v})})$ with the optimal contracts $\mathcal{C}^{\star}(b^{\mathrm{C}(\mathrm{v})})$ are well-defined and unique, following
$\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})=\theta^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{b^{\mathrm{w}}}(w)\;,$ (7)
as well as the best utility $\mathrm{v}$ can obtain from action $a$ and any couple $b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}$ as
$\mu^{\star,\mathrm{v}}(a)=\max_{b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}}\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})\;.$ (8)
We also have that
$\sum_{\mathrm{v}\in\mathrm{V}}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{\text{Card}(\mathrm{V})}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})=\max_{(a^{\mathrm{v}})\in\mathcal{A}^{\text{Card}(\mathrm{V})}}\sum_{\mathrm{v}\in\mathrm{V}}\theta(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})\;.$
The quantities arising in Lemma 4.1 are well-defined for the leafs, for which it is easy to quantify them (see (4)). Then, an induction allows to define these quantities for the upper layers (see the proofs in Appendix B). In the last equation from Lemma 4.1, the left hand side represents the target of the players playing on their own while the right hand side represents the target of collaborative players.
Equilibrium. Lemma 4.1 makes the contracts explicit and relates Equation 8 to a notion of equilibrium. Once the players in $\mathrm{V}$ take the set of actions maximizing $\sum_{\mathrm{v}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})$, the cost of deviating for any player is larger than the reward she can expect from a deviated action. This is why our optimal contracts are closely linked with a notion of equilibrium.
Regret Definition. Since we define $\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})$ as the utility for player $\mathrm{v}$ with action $a$ and her children playing actions $b^{\mathrm{C}(\mathrm{v})}$, we can now define her regret for the time window $\{s,\ldots,t\}$, implying the optimal contracts in hindsight as well as the received contract, following
$\mathfrak{R}^{\mathrm{v}}(s,t)$ $=\sum_{l=s}^{t}\max_{a,b^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})}\right\}$
$-\sum_{l=s}^{t}\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})$ (9)
$-\sum_{l=s}^{t}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{B_{l}^{\mathrm{w}}}(A^{\mathrm{w}}_{l})\tau_{l}(\mathrm{w})}\right\}.$
Here, we define formally what is called a pseudo-regret (see, Auer & Chiang, 2016; Lattimore & Szepesvári, 2020), since it is not written with an expectation for the whole sum and is a stronger notion than the expected regret. Also, this notion of regret for our game takes into account the contract $(B_{t}^{\mathrm{v}},\tau_{t}(\mathrm{v}))$ that player $\mathrm{v}$ receives from $\mathrm{P}(\mathrm{v})$ at each step $l\in\{s,\ldots,t\}$. This is coherent in the sense that if a huge transfer is offered to $\mathrm{v}$ for some specific action $B^{\mathrm{v}}$, it would be a bad move not to play it - hence the presence of $\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})$ in the regret. Finally, we define $\mathfrak{R}^{\mathrm{v}}$ on any window $\{s,\ldots,t\}\subseteq\mathbb{N}$ because analyzing regret on specific time intervals is crucial in this work. For the ease of notation, we also define the regret over the whole horizon as
$\mathfrak{R}^{\mathrm{v}}(T)=\mathfrak{R}^{\mathrm{v}}(1,T)\;,$
and we now introduce a key lemma that exposes the various sources of regret experienced by a player.
Lemma 4.2.
Considering steps $\{s,\ldots,t\}\subseteq\{1,\ldots,T\}$ and a player $\mathrm{v}\in\mathrm{V}$ who offers contracts $(B_{l}^{\mathrm{C}(\mathrm{v})},\tau_{l}(\mathrm{C}(\mathrm{v})))$ at step $l$ to her children, we can decompose her regret into several terms, following
$\mathfrak{R}^{\mathrm{v}}(s,t)\leqslant\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}(s,t)+\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}(s,t)+\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}(s,t)\;,$
where the action regret is defined as
$\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}(s,t)=$ $\sum_{l=s}^{t}\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{a}+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}$
$-\sum_{l=s}^{t}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v}))}\right\}\;,$
the payment regret is defined as
$\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}(s,t)=\sum_{l=s}^{t}\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\tau_{B_{l}^{\mathrm{w}}}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w}))_{+}\;,$
and the deviation regret is defined as
$\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}(s,t)=\sum_{l=s}^{t}(\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}))_{+}\;.$
Note that the action regret $\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}$ corresponds to player $\mathrm{v}$’s error in the action she chooses to play. The payment regret $\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}$ corresponds to the payment excess of $\mathrm{v}$ to her children with respect to the optimal transfers while the deviation regret $\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}$ corresponds to how much the children mis-responded to the contracts.
Good Behavior. The following assumption about the players’ action regret is the most decisive tool of this work. Intrinsically, our results hold because we are able to show that if the children of a node satisfy this assumption, then this node satisfies the assumption as well, with potentially different parameters. Similar assumptions are often done in the literature (see, e.g., Donahue et al., 2024) when online principal-agent interactions are considered.
Assumption 4.3.
Considering a node $\mathrm{v}\in\mathrm{V}$, there exists parameters $(\mathrm{W}_{\mathrm{v}},\mathrm{C}_{\mathrm{v}},\kappa_{\mathrm{v}},\zeta_{\mathrm{v}})$ such that for any steps $s>\mathrm{W}_{\mathrm{v}}$, $t\in\mathbb{N}^{\star}$ with $s+t\leqslant T$ and contracts from $\mathrm{P}(\mathrm{v})$ such that $(B_{l}^{\mathrm{v}},\tau_{l}(\mathrm{v}))=(B_{l}^{\mathrm{v}},\tau_{B_{l}^{\mathrm{v}}})$ for any step $l\in\{s+1,\ldots,s+t\}$, then
$\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}(s+1,s+t)\leqslant\mathrm{C}_{\mathrm{v}}\,t^{\kappa_{\mathrm{v}}}\;,$
with probability at least $1-1/t^{\zeta_{\mathrm{v}}}$. We call $\mathrm{W}_{\mathrm{v}}$ the time to learn associated with player $\mathrm{v}$.
4.3 is written for the Action Regret in order to avoid a dramatic propagation of the errors from the leafs to the roots, as explained in the penultimate paragraph of the section. The literature generally considers single principal-agent interactions (Haghtalab et al., 2022; Mansour et al., 2022). This work is one of the first in the learning community to study nested interactions between players, as well as how the corresponding parameters evolve along the tree. This is why it requires some new techniques, such as 4.3. Consider a leaf $\mathrm{v}\in\mathrm{V}$: there are no children, hence $\mathrm{C}(\mathrm{v})=\varnothing$ and $\mu^{\star,\mathrm{v}}(a)=\theta^{\mathrm{v}}(a)$ for any $a\in\mathcal{A}$. Thus, her action regret can be written
$\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}(s+1,s+t)$ $=\sum_{l=s+1}^{s+t}\max_{a\in\mathcal{A}}\{\theta(a)+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}$
$-\sum_{l=s+1}^{s+t}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v}))}\right\},$
and Scheid et al. (2024a, Proposition 2) show that in that case, if $\mathrm{v}$ uses a UCB subroutine (see Algorithm 3), 4.3 is satisfied with constants $(0,8\sqrt{K\log(KT^{3})},1/2,2)$. It also holds with other generic bandit algorithms, such as Explore Then Commit for instance.
Our goal is to design an algorithm that is no-regret for a player $\mathrm{v}$, when all of her children $\mathrm{w}$ satisfy 4.3. Additionally, we want the player $\mathrm{v}$ to also satisfy 4.3 when following this algorithm. In that purpose, we introduce the algorithm COBRA, whose pseudo-code is provided in Algorithm 1. We describe below COBRA more precisely for a player $\mathrm{v}$, assuming her children all satisfy 4.3.
Our Algorithm. It is impossible to learn from agents who do not behave properly. This is why the first step of COBRA is to wait $\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{W}_{\mathrm{w}}$ steps so the children best-respond almost all of the time to the contracts (for instance in Lin & Chen, 2024, the agent also approximately best-responds, which gives power to the principal). Since the children all satisfy 4.3, we are then ensured that they have a small regret. Thus, we use an epoched binary search-like procedure to get an upper estimate $(\hat{\tau}_{b}(\mathrm{w}))_{b\in\mathcal{A},\mathrm{w}\in\mathrm{C}(\mathrm{v})}$ of the optimal payments, following (11). This is achieved using the auxiliary procedure Exp Sub (see Algorithm 2).
Then, COBRA uses a black-box bandit subroutine (typically UCB, see, e.g., Algorithm 3) which runs on a $K^{\mathrm{B}+1}$-armed bandit instance with shifted rewards defined as $\theta^{\mathrm{v}}(A,B^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{B^{\mathrm{w}}}(\mathrm{w})$. If there was more structure on the rewards, a combinatorial bandit algorithm – e.g., CUCB – could be used instead. We call a black-box subroutine any bandit algorithm that runs on a multi-armed bandit instance and is no-regret. In our case, the arms’ mean rewards $\theta(a,b^{\mathrm{C}(\mathrm{v})})$ are shifted by the needed transfers $-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{b}(\mathrm{w})$. For a set of actions $(A,B^{\mathrm{C}(\mathrm{v})})$ recommended by the black-box subroutine, COBRA offers a contract $(B^{\mathrm{C}(\mathrm{v})},\hat{\tau}_{B^{\mathrm{C}(\mathrm{v})}}(\mathrm{C}(\mathrm{v})))$ to her children. Note that the leafs run the black-box subroutine from the start ($t=1$) since they do not have children and hence do not offer contracts.
Algorithm 1 COBRA (player $\mathrm{v}$).
  Input: Set of action $\mathcal{A}$, horizon $T$, subroutine $\pi^{\mathrm{w}}$ and parameters $\mathrm{C}_{\mathrm{w}}$, $\mathrm{W}_{\mathrm{w}}$, $\kappa_{\mathrm{w}}$, $\zeta_{\mathrm{w}}$ for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, parameters $\alpha,\beta,\eta$
  Compute $\mathcal{H}^{\mathrm{v}}_{0}=\varnothing,\mathrm{W}=\max_{\mathrm{w}}\mathrm{W}_{\mathrm{w}},\mathrm{C}=\max_{\mathrm{w}}\mathrm{C}_{\mathrm{w}},\kappa=\max_{\mathrm{w}}\kappa_{\mathrm{w}},\zeta=\min_{\mathrm{w}}\zeta_{\mathrm{w}}$
  Pick action $B\in\mathcal{A}$ at random.
  for $t=1,\ldots,\mathrm{W}$ do
     Offer contract $(B,0)$ to any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$
     Play action $A_{t}^{\mathrm{v}}\sim\mathrm{Unif}(K)$.
  end for
  for any agent $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ do
     for any action $b\in\mathcal{A}$ do
        $\underline{\tau}^{\mathrm{w}},\bar{\tau}^{\mathrm{w}}=\texttt{Exp Sub}(T,\mathrm{w},\pi^{\mathrm{w}},b,\mathrm{W},\mathrm{C},\kappa,\alpha,\beta,\eta)$
        Compute $\hat{\tau}_{b}(\mathrm{w})=\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)+1/T^{\beta}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}$
        # See Algorithm 2
     end for
  end for
  for $t=\mathrm{W}+K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil+1,\ldots,T$ do
     Get set of actions from the $K^{\mathrm{B}+1}$ bandit instance: $(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{w})})=\texttt{Bandit Sub}(U_{t},\mathcal{H}_{t-1}^{\mathrm{v}},B^{\mathrm{v}}_{t},\tau_{t}(\mathrm{v}))$.
     Play action $A_{t}^{\mathrm{v}}$, offer contract $\mathcal{C}_{t}^{\mathrm{w}}=(B_{t}^{\mathrm{w}},\hat{\tau}_{B_{t}^{\mathrm{w}}}(\mathrm{w}))$ to any player $\mathrm{w}\in\mathrm{C}(\mathrm{v})$.
     Observe $A_{t}^{\mathrm{w}}=\pi^{\mathrm{w}}(U_{t}^{\mathrm{w}},\mathcal{H}_{t-1}^{\mathrm{w}},B_{t}^{\mathrm{w}},\tau_{t}(\mathrm{w}))$ for any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$
     Update history $\mathcal{H}_{t}^{\mathrm{v}}$ and history $\mathcal{H}_{t}^{\mathrm{w}}$ for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$
  end for
4.3 ensures that the children actions $(A^{\mathrm{C}(\mathrm{v})})$ do not deviate too much from the recommended action $(B^{\mathrm{C}(\mathrm{v})})$. With mean-based algorithms such as UCB, the contract $(B^{\mathrm{v}}_{t},\tau_{t}(\mathrm{v}))$ received from $\mathrm{P}(\mathrm{v})$ at each step $t$ can simply be added to the score of each arm to be taken into account. Based on this assumption, Lemma 4.4 shows that if $\mathrm{v}$ uses the algorithm COBRA and all her children satisfy 4.3, $\mathrm{v}$ actually satisfies the assumption in turn.
Once the leafs satisfy 4.3, their parents – if they use COBRA – also satisfy 4.3 with new constants, again by Lemma 4.4. By induction, 4.3 is then satisfied by all the nodes up to the root if everyone plays COBRA, where the convergence rates are detailed in Lemma 4.5. In that way, we show that if all the players use COBRA, then they converge to the optimum as a whole.
Before moving to the results, we first describe more precisely the different subroutines used in COBRA. Assuming that the children $\mathrm{C}(\mathrm{v})$ of a node $\mathrm{v}$ satisfy 4.3 with parameters $(\mathrm{C}_{\mathrm{w}},\mathrm{W}_{\mathrm{w}},\kappa_{\mathrm{w}},\zeta_{\mathrm{w}})$, we can define
$\mathrm{C}=\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{C}_{\mathrm{w}}\;\;,\;\;\mathrm{W}=\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{W}_{\mathrm{w}}\;,$ (10)
$\kappa=\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\kappa_{\mathrm{w}}\;\;,\;\;\zeta=\min_{\mathrm{v}\in\mathrm{C}(\mathrm{w})}\zeta_{\mathrm{w}}\;.$
Contract Exploration Subroutine. Algorithm 1 is based on the payment exploration subroutine Exp Sub which estimates $\tau^{\star}_{b}(\mathrm{w})$ in the range $[\underline{\tau}^{\mathrm{w}}_{b},\bar{\tau}^{\mathrm{w}}_{b}]$ and shrinks the gap $\bar{\tau}^{\mathrm{w}}_{b}(\lambda)-\underline{\tau}^{\mathrm{w}}_{b}(\lambda)$ along the iterations $\lambda$. More precisely, consider any player $\mathrm{v}$ and action $b\in\mathcal{A}$. Thanks to 3.1, Exp Sub starts with an estimate $[\underline{\tau}^{\mathrm{w}}_{b}(0),\bar{\tau}^{\mathrm{w}}_{b}(0)]=[0,1]$ that contains $\tau^{\star}_{b}(\mathrm{w})$ almost surely. For a length-$\lceil T^{\alpha}\rceil$ batch $\lambda$ of iterations, $\mathrm{v}$ offers a contract $(b,(\bar{\tau}^{\mathrm{w}}_{b}(\lambda)+\underline{\tau}^{\mathrm{w}}_{b}(\lambda))/2)$ to $\mathrm{w}$. At the end of the batch $\lambda$, the number of steps for which the contract has been accepted allows to guess whether $\tau^{\star}_{b}(\mathrm{w})>(\bar{\tau}^{\mathrm{w}}_{b}(\lambda)+\underline{\tau}^{\mathrm{w}}_{b}(\lambda))/2$ or $\tau^{\star}_{b}(\mathrm{w})<(\bar{\tau}^{\mathrm{w}}_{b}(\lambda)+\underline{\tau}^{\mathrm{w}}_{b}(\lambda))/2$. Exp Sub iterates until a precision $1/T^{\beta}$ is obtained for $\tau^{\star}_{b}(\mathrm{w})$, after $\lceil\log T^{\beta}\rceil$ iterations.
When player $\mathrm{v}$ runs Algorithm 2 with parameters $\alpha,\beta,\eta$, at the end of the $K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil$ steps of the payment exploration phase, we can define her estimated transfer $\hat{\tau}_{b}(\mathrm{w})$ for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ and any action $b\in\mathcal{A}$, following
$\hat{\tau}_{b}(\mathrm{w})=\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)+1/T^{\beta}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}\;.$ (11)
The quantity $1/T^{\beta}$ ensures that $\hat{\tau}_{b}(\mathrm{w})$ is above $\tau^{\star}_{b}(\mathrm{w})$ with high probability while the term $\mathrm{C}\,\mathrm{B}\,T^{-\eta}$ is there to make action $b$ really better then the others and hence, force $\mathrm{w}$ to play it. As it is shown formally in Proposition B.5 from Appendix B, any player $\mathrm{v}\in\mathrm{V}$ running Exp Sub with her children satisfying 4.3, obtains an upper bound $\hat{\tau}_{b}(\mathrm{w})$ of the optimal payment such that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A}$, we have
$|\hat{\tau}_{b}(\mathrm{w})-\tau^{\star}_{b}(\mathrm{w})|=o(T)\;.$ (12)
After the payment exploration phase, suppose that a player $\mathrm{v}$ takes action $A_{t}^{\mathrm{v}}$ at step $t\in[T]$ and offers a contract $(B_{t}^{\mathrm{C}(\mathrm{v})},\hat{\tau}_{B_{t}^{\mathrm{C}(\mathrm{v})}}(\mathrm{C}(\mathrm{v})))$ to her children. If they all best-respond, then $(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})=(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})$, which means that we can define the quantity $Z_{t}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})$ that is the reward observed by player $\mathrm{v}$, taking into account the cost of the actions, following
$Z_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})=X_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{\mathrm{w}}(B_{t}^{\mathrm{w}})\;,$
as well as the true mean of this quantity
$\nu_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})=\theta_{t}^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{\mathrm{w}}(B_{t}^{\mathrm{w}})\;.$
We can now state our most important technical lemma, which allows the nodes of the tree to play well enough as we go from the leafs to the nodes. If node $\mathrm{v}$ runs COBRA with parameters $(\eta,\alpha,\beta)$ and her children satisfy 4.3 with parameters $(\mathrm{W},\mathrm{C},\kappa,\zeta)$, we need to have
$\beta/\alpha<1-\kappa\;,$ (13)
for the payment exploration subroutine to be successful (see Appendix B, Lemma B.1 for more details). As shown in (11), $\mathrm{B}\mathrm{C}T^{-\eta}$ is the amount of extra payment added to each arm to ensure that the children accept the contracts for enough steps. The regret bounds imply both a term evolving as $\mathcal{O}(T^{\kappa+\eta})$ and a term evolving as $\mathcal{O}(T^{1-\eta}$, hence the need of choosing $\eta$ such that $0<\eta<1-\kappa$, which is satisfied with our choice of parameters in (14).
Lemma 4.4.
Consider a player $\mathrm{v}\in\mathrm{V}$ running COBRA with parameters $(\eta,\alpha,\beta)$ such that any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ satisfies 4.3 with parameters $(\mathrm{W},\mathrm{C},\kappa,\zeta)$ and $\beta/\alpha<1-\kappa$. Assume that 3.1 holds. In that case, $\mathrm{v}$ also satisfies 4.3 with parameters $(\mathrm{W}+K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,10\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3})},\max\{1/2,\kappa+\eta,1-\beta\},\tilde{\zeta})$ with $\tilde{\zeta}=\alpha\zeta-\log(4\mathrm{B}K\log T)/\log T$.
Tree Structure. We now specifically study the game within the tree structure $\mathcal{T}$. Depending on the depth $d\in\{2,\ldots,D\}$, the parameters $(\eta_{d},\alpha_{d},\beta_{d})$ are fine-tuned. In our setting for the next theorems, $(\mathrm{W}_{d},\mathrm{C}_{d},\kappa_{d},\zeta_{d},\alpha_{d},\beta_{d})$ are the layers’ parameters and are common to any node $\mathrm{v}$ that is located at depth $d$ in the tree. Assume that for any $d\in\{2,\ldots,D\}$, a node $\mathrm{v}$ located at depth $d$ runs COBRA with parameters
$\eta_{d}=\frac{1}{2d(d-1)}\;,\;\alpha_{d}=\frac{(D+1)(d-1)}{D\,d}\;,\;\beta_{d}=\frac{1}{2d}\;,$ (14)
where $(\alpha,\beta,\eta)$ are not defined for the leafs which run directly Algorithm 3 without Algorithm 2. Lemma 4.4 is the key ingredient to prove our regret bounds. An important feature is the extra payment $\mathrm{C}\,\mathrm{B}\,T^{-\eta}$ in $\hat{\tau}$, which creates a larger reward gap and forces the children to play the action recommended by the parent. While this extra payment largely increases the payment regret, it yields a better control of the deviation regret as it makes deviating from the recommended actions more costly for the children. Moreover, any children deviation yields feedback on different arms than the ones intended by the principal and thus leads to lost exploration steps. As a consequence, this extra payment is not only needed to control the deviation regret, but also the action regret.
Lemma 4.5.
Consider a large enough horizon so that $\log T\geqslant D^{2}\log(4\mathrm{B}K\log T)$. Assume that any player at depth $d$ in the tree runs COBRA with the parameters defined in (14). Then, any player at depth $d\in\{1,\ldots,D\}$ satisfies 4.3 with parameters $(d\,K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,\mathrm{C},1-1/2d,\zeta_{d})$, and we have $\zeta_{d}\geqslant 1/d$.
The assumption $\log T\geqslant D^{2}\log(4\mathrm{B}K\log T)$ may seem strong. However, it only reflects the fact that the horizon needs to be large enough as compared to the size of the tree, which is natural since the opposite would not allow players to learn from others within the available amount of time.
Theorem 4.6.
Consider the game running for a horizon $T$, large enough so that $\log T\geqslant D^{2}\log(4\mathrm{B}K\log T)$. Assume that any player in the tree runs Algorithm 1 with the parameters specified in (14). In that case, for any player $\mathrm{v}\in\mathrm{V}$ at depth $d$ in the tree, we have that
$\mathfrak{R}^{\mathrm{v}}(T)=\tilde{\mathcal{O}}(T^{1-\frac{1}{2d^{2}}})\;,$
with probability at least $1-T^{\zeta_{d}}$ for $\zeta_{d}\geqslant 1/d$, and $\tilde{\mathcal{O}}$ hiding logarithmic factors.
In particular, Theorem 4.6 implies that we have
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{v}}(T)}\right]=o(T)\;,$
and hence COBRA is actually a no-regret strategy for the players in the tree.
This result is exciting because the orchestration of the whole tree appears without the need of any kind of collaboration nor centralized leader. The single fact of allowing contracts makes the players behave optimally and maximize the social welfare, as it is formulated clearly in Corollary 4.7, from a coalition point of view. This result shows that, although being selfish on their own, the players can converge together towards the global equilibrium. This is accomplished through repeated contracts between the players, which helps avoiding globally suboptimal actions at any step. Note that the scaling of the regret depends on your depth in the tree: you have a lower regret when you interact with better-responding agents - i.e. lower ranked in the depth.
Corollary 4.7.
Suppose that the horizon is large enough so that $\log T\geqslant D^{2}\log(4\mathrm{B}K\log T)$ and that any player $\mathrm{v}\in\mathrm{V}$ runs Algorithm 1 with the parameters specified in (14). In that case, we have that
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right]\\ =o(T)\;.$ (15)
Why Decomposing the Regret? 4.3 could be stated in terms of the overall regret $\mathfrak{R}^{\mathrm{v}}$ and it would still yield convergence guarantees. However, the resulting regret bound would scale as $T^{1-1/2^{D}}$, which is significantly worse than the $T^{1-1/D^{2}}$ obtained through our decomposition. Beyond enhancing clarity, decomposing the regret as we did in Lemma 4.2 is crucial to mitigate the dramatic propagation of regret terms.
Speed of convergence. An interesting feature is the evolution of the regret’s speed of convergence along the layers. We observe an importance decay of the upper bounds as we descent down to the leafs - the worst regret bound being for the root. It can be understood as follows: the better your agent behaves, the lower is your regret - the optimal scenario being contracting with a best-responding agent. Here, we showed that contracting is still possible among a tree-structure of players but we still observe a propagation of the errors from an agent to his principal.
On the necessity of transfers. One could ask why the transfers were introduced in this sequential multi-agent game. As shown in Scheid et al. (2024a) in the simpler two-players case, without transfers, one would necessarily have
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})\geqslant\Delta T\;,$
for some reward gap $\Delta>0$. It is directly linked to the fact that without transfers, the utilities of the players are not aligned, which creates inefficiencies and results in a loss of social welfare. Without contracting in this multi-agent game, the global utility breaks down. Hence, contracting restores the efficiency of the game.
5 Conclusion
We showed that allowing contracts between players allows non-collaborative no-regret learners to converge to the global optimum within an intricate principal-agent tree structure. Algorithmically, the first step for the players is to learn the optimal contract needed to convince their agents to play some specific action. Then, a bandit subroutine can be used on the shifted instance. Interestingly, the players end up behaving as if they were collaborating. The straightforward extensions of this work would be similar studies with dynamic trees, additional context or in a Markov Decision Process scenario. It would also be interesting to relax to some extent the myopic and regret-maximizing assumptions on our agents. Finally, there are a lot of applications of our setup in auctions, collaborative games or specific mechanism design questions, since we tried to provide a general setup to encompass various questions.
Acknowledgements
Funded by the European Union (ERC, Ocean, 101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Detailed Explanations on the Algorithms
Payment Exploration Phase for Player $\mathrm{v}\in\mathrm{V}$. A first step is to explain how works the payment exploration subroutine Algorithm 2, which roughly corresponds to a batched binary search operated by the parent node $\mathrm{v}\in\mathrm{V}$ on all of her children $\mathrm{w}\in\mathrm{C}(\mathrm{v})$. As presented in Lemma 4.1, the true rewards (which takes into account the transfers to the children) of the actions can be defined by induction. The goal of a player $\mathrm{v}\in\mathrm{V}$ is to learn first the optimal transfer $\tau^{\star}_{b}(\mathrm{w})$ for any $b\in\mathcal{A},\mathrm{w}\in\mathrm{C}(\mathrm{v})$ which then gives the optimal contract $(b,\tau^{\star}_{b}(\mathrm{w}))$. For that purpose, $\mathrm{v}$ first waits $\mathrm{W}=\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{W}_{\mathrm{w}}$ steps, so that all her children almost best-respond on their own bandit instances with their own children nodes. After these steps, the player $\mathrm{v}$’s payment exploration phase starts. The binary search strategy is the same across all the children. More precisely, for any $b\in\mathcal{A}$, we define
$\Psi_{b}^{\mathrm{v}}=\mathrm{W}+(b-1)\,\lceil\log T^{\beta}\rceil\,\lceil T^{\alpha}\rceil\;,$ (16)
which is the step at which $\mathrm{v}$ starts the binary search procedure to estimate arm $b$’s associated optimal transfer $\tau^{\star}_{b}(\mathrm{w})$ for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$. For any binary search step $d\in\{1,\ldots,\lceil\log T^{\beta}\rceil\}$, we also define
$k_{b,d}^{\mathrm{v}}=\Psi_{b}^{\mathrm{v}}+(d-1)\,\lceil T^{\alpha}\rceil\;,$ (17)
which corresponds to the step after which $\mathrm{v}$ starts the $d$-th batch of binary search on her children’s arm $b$. Now, we can define
[…]
which corresponds to the number of steps during the payment exploration phase on arm $b$, player $\mathrm{w}$ when the child did not accept the transfer. We also define our ”nice” event $\mathcal{E}^{\mathrm{w}}_{\text{good}}$ for any player $\mathrm{w}\in\mathrm{V}$ on a window of steps $[s+1,\ldots,s+t]$ as
$\mathcal{E}^{\mathrm{w}}_{\text{good}}([s+1,\ldots,s+t],\mathrm{C},\kappa)=\mathopen{}\mathclose{{}\left\{\sum_{l=s+1}^{s+t}\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{w}}_{a}+\text{1}_{B_{l}^{\mathrm{w}}}(a)\,\tau_{l}(\mathrm{w})\}-\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{w}}(A_{l}^{\mathrm{w}},B_{l}^{\mathrm{C}(\mathrm{w})})+\text{1}_{B_{l}^{\mathrm{w}}}(A_{l}^{\mathrm{w}})\,\tau_{l}(\mathrm{w})}\right\}\leqslant\mathrm{C}t^{\kappa}}\right\}\;,$ (19)
and any node $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ which satisfies 4.3 with parameters $(\mathrm{W},\mathrm{C},\kappa,\zeta)$, also satisfies $\mathcal{E}^{\mathrm{w}}_{\text{good}}([s+1,\ldots,s+t],\mathrm{C},\kappa)$ for any $s>\mathrm{W}$.
Algorithm 2 Exp Sub (player $\mathrm{v}$).
  Input: $T,\mathrm{w},\pi^{\mathrm{w}},b,\mathrm{W},\mathrm{C},\kappa,\alpha,\beta,\eta$.
  Initialize $\underline{\tau}^{\mathrm{w}}_{b}(0)=0,\bar{\tau}^{\mathrm{w}}_{b}(0)=1,\tilde{T}=\lceil T^{\alpha}\rceil$
  for $\lambda=0,\ldots,\lceil\log(T^{\beta})\rceil-1$ do
[…]
     for $t=\lambda\tilde{T}+1,\ldots,\lambda\tilde{T}+\tilde{T}$ do
        Propose contract $(b,\tau^{\text{mid},\mathrm{w}}_{b}(\lambda))$ to player $\mathrm{w}$
        $A^{\mathrm{w}}_{t}=\pi^{\mathrm{w}}(\mathcal{U}_{t+1}^{\mathrm{w}},\mathcal{H}_{t}^{\mathrm{w}},b,\tau^{\text{mid},\mathrm{w}}_{b}(\lambda))$
        if $A_{t}^{\mathrm{w}}\neq b$then
[…]
        end if
        Update player’s $\mathrm{w}$ history $\mathcal{H}^{\mathrm{w}}_{t+1}$ following the definition
     end for
[…]
        Return $\underline{\tau}^{\mathrm{w}}_{b}(\lambda),\bar{\tau}^{\mathrm{w}}_{b}(\lambda)$.
[…]
        $\bar{\tau}^{\mathrm{w}}_{b}(\lambda)=\tau^{\text{mid},\mathrm{w}}_{b}(\lambda)+1/T^{\beta}$ and update history $\mathcal{H}^{\mathrm{v}}_{t}$.
     else
        $\;\underline{\tau}^{\mathrm{w}}_{b}(\lambda)=\tau^{\text{mid},\mathrm{w}}_{b}(\lambda)-1/T^{\beta}$ and update history $\mathcal{H}^{\mathrm{v}}_{t}$.
     end if
  end for
UCB Subroutine. We present the player’s UCB subroutine. As explained above, once the optimal contracts are estimated by Algorithm 2, COBRA uses Algorithm 3 to be no-regret on her shifted bandit instance (the reward of each arm is shifted by the contract that needs to be paid so the children actually play the right action). As shown in Scheid et al. (2024a), the policy Algorithm 3 satisfies 4.3 with $\kappa=1/2,\zeta=2$. We make use of this for the proofs in Appendix B. Since the leafs of the tree directly use Algorithm 3 without the need of offering contracts, they are guaranteed to satisfy 4.3.
Algorithm 3 UCB’s subroutine - Player $\mathrm{v}$
1:  Input: Set of arms $\mathcal{A}$, horizon $T$, $\eta$, estimated incentives $(\hat{\tau}_{b}(\mathrm{w}))_{\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A}}$
2:  Initialize: For any combination of arms $(a,b^{\mathrm{C}(\mathrm{v})})\in\mathcal{A}^{\mathrm{B}+1}$, set $\hat{\nu}_{a,b^{\mathrm{C}(\mathrm{v})}}=0,\,T_{a}=0$
3:  for $1\leqslant t\leqslant K^{\mathrm{B}+1}$do
4:     Pull an arm $A_{t},B_{t}^{\mathrm{C}(\mathrm{v})}$ that has not been pulled yet.
5:     Update $\hat{\nu}_{A_{t},B_{t}^{\mathrm{C}(\mathrm{v})}}=X_{A_{t},B_{t}^{\mathrm{C}(\mathrm{v})}}(t)-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{B_{t}^{\mathrm{w}}}(\mathrm{w}),\,T_{A_{t},B_{t}^{\mathrm{C}(\mathrm{v})}}(t)=1$
6:  end for
7:  for $t\geqslant K^{\mathrm{B}+1}+1$ do
8:     Observe the contract $(B_{t}^{\mathrm{v}},\tau_{t}(\mathrm{v}))$.
9:     Choose actions $(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})\in\operatorname*{\arg\!\max}_{a,b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}+1}}\mathopen{}\mathclose{{}\left\{\hat{\nu}_{a,b^{\mathrm{C}(\mathrm{v})}}(t-1)+2\sqrt{\frac{\log(K^{\mathrm{B}+1}\,T^{3})}{T_{a,b^{\mathrm{C}(\mathrm{v})}}(t-1)}}+\text{1}_{B_{t}^{\mathrm{v}}}(a)\tau_{t}(\mathrm{v})}\right\}$
10:     if $A_{t}^{\mathrm{C}(\mathrm{v})}=B_{t}^{\mathrm{C}(\mathrm{v})}$ then
11:        Update $T_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t)=T_{\mathcal{A}_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t-1)+1,\,\hat{\nu}_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t)=\frac{1}{T_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t)}(T_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t-1)\hat{\nu}_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t-1)+X_{A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})}}(t)-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{B_{t}^{\mathrm{w}}}(\mathrm{w}))$
12:     end if
13:  end for
Appendix B Proofs
Proof of Lemma 4.1.
We conduct the proof by induction starting from the first layer $d=1$ contracting with the leafs and ascending to the root.
Initialization. The optimal utility any leaf $\mathrm{v}$ can extract from the arms is $\theta^{\mathrm{v}}(a)$ for any action $a\in\mathcal{A}$. Therefore, we have that for $\mathrm{v}$, the average optimal utility is defined as
$\mu^{\mathrm{v}}(a)=\theta^{\mathrm{v}}(a)\;\text{ and }\;\mu^{\star,\mathrm{v}}(a)=\theta^{\mathrm{v}}(a)\;,$
since $\mathrm{C}(\mathrm{v})=\varnothing$ and hence $\mathrm{v}$’s contracts are empty. As we show it in the main text, we also have that for any node $\mathrm{u}$ in the first layer (in interaction with children $\mathrm{v}\in\mathrm{C}(\mathrm{u})$ being leafs), we have that
$\tau^{\star}_{b^{\mathrm{v}}}(\mathrm{v})=\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{v}}(b)-\mu^{\star,\mathrm{v}}(b^{\mathrm{v}})=\max_{b\in\mathcal{A}}\theta^{\mathrm{v}}(b)-\theta^{\mathrm{v}}(b^{\mathrm{v}})\;,$
and this quantity is well-defined and unique. Considering $\mathrm{u}=\mathrm{P}(\mathrm{v})$, it leads to
$\mu^{\mathrm{u}}(a,b^{\mathrm{C}(\mathrm{v})})=\theta^{\mathrm{u}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{v}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})\;,$
which is well defined too. Finally, for any $a\in\mathcal{A},\mu^{\star,\mathrm{u}}(a)$ is well-defined and unique as a maximum over a finite number of scalars. Therefore, our initialization holds.
Induction step. Consider a player $\mathrm{v}\in\mathrm{V}$ that is not a leaf. We consider a single child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ to show our result. Our induction assumption gives us that $\mu^{\star,\mathrm{w}}(b)$ is well-defined for any $b\in\mathcal{A}$. For any $b^{\mathrm{w}}\in\mathcal{A}$, picking $b^{\mathrm{w}}$ can lead in hindsight to an optimal average utility of $\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})$ by definition. Since $\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})$ is defined as the smallest transfer such that $b^{\mathrm{w}}$ is the best action $\mathrm{w}$ could play in hindsight, we necessarily have that for any $b\in\mathcal{A}$
$\text{1}_{b^{\mathrm{w}}}(b^{\mathrm{w}})\,\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})+\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})\geqslant\text{1}_{b^{\mathrm{w}}}(b)\,\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})+\mu^{\star,\mathrm{w}}(b)\;,$ (20)
which gives that necessarily
$\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})\geqslant\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})\;.$
As long as $\tau_{b^{\mathrm{w}}}(\mathrm{w})>\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})$, (20) is satisfied with a strict inequality: which makes $b^{\mathrm{w}}$ the best possible action in hindsight. Hence, taking the infimum of $\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})$ gives that
$\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})=\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(b^{\mathrm{w}})\;,$
hence the first part of the result. For any couple $(a,b^{\mathrm{C}(\mathrm{w})})\in\mathcal{A}\times\mathcal{A}^{\mathrm{B}}$, if the players $\mathrm{C}(\mathrm{w})$ are fully rational and have a full knowledge of the game, then offering to any player $\mathrm{w}\in\mathrm{C}(\mathrm{w})$ the contract $\mathcal{C}^{\mathrm{w}}=(b^{\mathrm{w}},\tau_{b^{\mathrm{w}}}(\mathrm{w}))$ such that $\tau_{b^{\mathrm{w}}}(\mathrm{w})\geqslant\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})$ ensures that $A^{\mathrm{w}}=b^{\mathrm{w}}$. This leads to the average utility
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{U}^{\mathrm{v}}(a,\mathcal{C}^{\mathrm{C}(\mathrm{v})})}\right]=\theta^{\mathrm{w}}(a,A^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{b^{\mathrm{w}}}(A^{\mathrm{w}})\,\tau_{b^{\mathrm{w}}}(\mathrm{w})=\theta^{\mathrm{w}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau_{b^{\mathrm{w}}}(\mathrm{w})\;,$
and finally taking the maximum over the transfers that only need to satisfy $\tau_{b^{\mathrm{w}}}(\mathrm{w})\geqslant\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})$ gives that
$\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})=\theta^{\mathrm{w}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})\;.$
Since $(\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})}))_{a\in\mathcal{A},b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}}$ are finite, unique and well-defined quantities, we have that
$\mu^{\star,\mathrm{v}}(a)=\max_{b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}^{\mathrm{B}}}\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})$
is also well-defined and unique. Therefore, our induction holds, hence the result.
For the second part of the proof, we first write $\mathrm{V}_{d}$ for the set of nodes $\mathrm{v}\in\mathrm{V}$ such that $\mathrm{v}$ is located at depth $d$ and we consider the quantity
$\sum_{\mathrm{v}\in\mathrm{V}}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{\text{Card}(\mathrm{V})}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})$ $=\sum_{\mathrm{v}\in\mathrm{V}_{1}}\max_{a^{\mathrm{v}}}\theta^{\mathrm{v}}(a^{\mathrm{v}})+\sum_{\mathrm{v}\in\mathrm{V}_{2}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{a^{\mathrm{w}}}(\mathrm{w})}\right\}+\ldots$
$\quad+\sum_{\mathrm{v}\in\mathrm{V}_{D}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{a^{\mathrm{w}}}(\mathrm{w})}\right\}$
$=\sum_{\mathrm{v}\in\mathrm{V}_{1}}\max_{a^{\mathrm{v}}}\theta^{\mathrm{v}}(a^{\mathrm{v}})+\sum_{\mathrm{v}\in\mathrm{V}_{2}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(a^{\mathrm{w}}))}\right\}+\ldots$
$\quad+\sum_{\mathrm{v}\in\mathrm{V}_{D}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(a^{\mathrm{w}}))}\right\}$
$=\sum_{\mathrm{v}\in\mathrm{V}_{1}}\max_{a^{\mathrm{v}}}\theta^{\mathrm{v}}(a^{\mathrm{v}})-\sum_{\mathrm{v}\in\mathrm{V}_{1}}\max_{a^{\mathrm{v}}}\theta^{\mathrm{v}}(a^{\mathrm{v}})+\sum_{\mathrm{v}\in\mathrm{V}_{2}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})+\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\theta^{\mathrm{w}}(a^{\mathrm{w}})}\right\}+\ldots$
$\quad+\sum_{\mathrm{v}\in\mathrm{V}_{D}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(a^{\mathrm{w}}))}\right\}$
$=\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}_{1}\cup\mathrm{V}_{2}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})}\right\}+\ldots$
$\quad+\sum_{\mathrm{v}\in\mathrm{V}_{D}}\max_{a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,a^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\max_{b\in\mathcal{A}}\mu^{\star,\mathrm{w}}(b)-\mu^{\star,\mathrm{w}}(a^{\mathrm{w}}))}\right\}\;,$
where by convention $\mathrm{C}(\mathrm{v})=\varnothing$ for any $\mathrm{v}\in\mathrm{V}_{1}$. Applying the same technique iteratively finally gives that
$\sum_{\mathrm{v}\in\mathrm{V}}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{\text{Card}(\mathrm{V})}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})=\max_{(a^{\mathrm{v}})\in\mathcal{A}^{\text{Card}(\mathrm{V})}}\sum_{\mathrm{v}\in\mathrm{V}}\theta(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})\;.$
Recall our main assumption 4.3 that is fundamental for the convergence of the algorithms. A large part of the proof deals with this assumption and how it can be satisfied.
Lemma B.1.
[…]
Proof of Lemma B.1.
The whole proof is done conditionally on the event $\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil],\mathrm{C}\tilde{T}^{\kappa})$, which holds with probability at least $1-T^{-\alpha\zeta}$, since $\mathrm{w}$ satisfies 4.3 with parameters $(\mathrm{W},\mathrm{C},\kappa,\zeta)$. For the ease of notation, we define $\tau_{b}=\tau^{\text{mid},\mathrm{w}}_{b}(\mathrm{w})$. First suppose that $\tau_{b}\geqslant\tau^{\star}_{b}(\mathrm{w})+1/T^{\beta}$. By definition of the optimal transfer, we have that
$\text{1}_{b}(b)\tau_{b}+\mu^{\star,\mathrm{w}}(b)$ $\geqslant\tau^{\star}_{b}(\mathrm{w})+\mu^{\star,\mathrm{w}}(b)+1/T^{\beta}$
$=\max_{b^{\prime}\in\mathcal{A}}\{\mu^{\star,\mathrm{w}}(b^{\prime})+\text{1}_{b}(b^{\prime})\tau_{b}\}+1/T^{\beta}$
$\geqslant\mu^{\mathrm{v}}(b^{\prime},B^{\mathrm{C}(\mathrm{w})})+\text{1}_{b}(b^{\prime})\tau_{b}+1/T^{\beta}\;,$
for any $b^{\prime},B^{\mathrm{C}(\mathrm{w})}\in\mathcal{A}\times\mathcal{A}^{\mathrm{B}}$. Hence $b$ is the best arm $\mathrm{w}$ could play during $\mathrm{v}$’s payment exploration batch $\{k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}\}$. Since $\mathrm{w}$ satisfies 4.3 with parameters $(\mathrm{W},\mathrm{C},\kappa,\zeta)$, we have that
[…]
Lemma B.2.
Assume that we run Algorithm 2 and consider some payment exploration batch $d\in\lceil\log T^{\beta}\rceil$ run on arm $b\in\mathcal{A}$. Then $0\leqslant\underline{\tau}^{\mathrm{w}}_{b}(d)\leqslant\tau^{\text{mid},\mathrm{w}}_{b}(d)\leqslant\bar{\tau}^{\mathrm{w}}_{b}(d)\leqslant 1$.
Proof of Lemma B.2.
We prove the result by induction. Considering some action $b\in\mathcal{A}$, the initialisation of the proof holds thanks to the initialisation of our algorithm: $\underline{\tau}^{\mathrm{w}}_{a}(0)=0,\bar{\tau}^{\mathrm{w}}_{a}(0)=1$ and therefore $\tau^{\text{mid},\mathrm{w}}_{a}(0)\in[\underline{\tau}^{\mathrm{w}}_{a}(0),\bar{\tau}^{\mathrm{w}}_{a}(0)]$. We now consider that $d$ payment exploration batches have been run on $b$. Suppose that we run an additional batch on action $a$. We have
$\tau^{\text{mid},\mathrm{w}}_{b}(d+1)=\frac{\bar{\tau}^{\mathrm{w}}_{b}(d)+\underline{\tau}^{\mathrm{w}}_{b}(d)}{2}\text{ which gives }\tau^{\text{mid},\mathrm{w}}_{b}(d+1)\in[\underline{\tau}^{\mathrm{w}}_{b}(d),\bar{\tau}^{\mathrm{w}}_{b}(d)]\;.$ (21)
After this iteration of binary search, we either have $\bar{\tau}^{\mathrm{w}}_{b}(d+1)=\tau^{\text{mid},\mathrm{w}}_{b}(d+1)+1/T^{\beta}$ and $\underline{\tau}^{\mathrm{w}}_{b}(d+1)=\underline{\tau}^{\mathrm{w}}_{b}(d)$ or $\bar{\tau}^{\mathrm{w}}_{b}(d+1)=\tau^{\text{mid},\mathrm{w}}_{b}(d)$ and $\underline{\tau}^{\mathrm{w}}_{b}(d+1)=\tau^{\text{mid},\mathrm{w}}_{b}(d+1)-1/T^{\beta}$. Therefore, we still have $0\leqslant\underline{\tau}^{\mathrm{w}}_{b}(d+1)\leqslant\tau^{\text{mid},\mathrm{w}}_{b}(d+1)\leqslant\bar{\tau}^{\mathrm{w}}_{b}(d+1)\leqslant 1$, hence the result for any $d\in\lceil\log T^{\beta}\rceil$ by induction. ∎
Lemma B.3.
Consider a player $\mathrm{v}\in\mathrm{V}$ who runs $D$ batches of binary search on action $b\in\mathcal{A}$ for all her children $\mathrm{C}(\mathrm{v})$. Assume that any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ satisfies 4.3 with parameters $(\mathrm{W}_{\mathrm{w}},\mathrm{C}_{\mathrm{w}},\kappa_{\mathrm{w}},\zeta_{\mathrm{w}})$ such that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, we have $\mathrm{W}_{\mathrm{w}}\leqslant\mathrm{W},\mathrm{C}_{\mathrm{w}}\leqslant\mathrm{W},\kappa_{\mathrm{w}}\leqslant\mathrm{W},\zeta_{\mathrm{w}}\geqslant\zeta$. Then, we have that
$\mathbb{P}\mathopen{}\mathclose{{}\left(\bigcap_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\bigcap_{d\in[D]}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil],\mathrm{C},\kappa)}\right)\geqslant 1-\mathrm{B}\,D/T^{\alpha\zeta}\;,$
which can be rewritten as: the following event
$\text{For any }\mathrm{w}\in\mathrm{C}(\mathrm{v}),d\in[D],b\in\mathcal{A},\sum_{l=k_{b,d}^{\mathrm{v}}+1}^{k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil}\mathopen{}\mathclose{{}\left\{\max_{b\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{b}+\text{1}_{B_{t}^{\mathrm{v}}}(b)\,\tau_{t}(\mathrm{v})\}-\mathopen{}\mathclose{{}\left(\mu^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\,\tau_{t}(\mathrm{v})}\right)}\right\}\leqslant\mathrm{C}\lceil T^{\alpha}\rceil^{\kappa},$
holds with probability at least $1-\mathrm{B}\,D/T^{\alpha\zeta}$.
Proof of Lemma B.3.
We do the proof considering the opposite event and applying an union bound. We have that
$\mathbb{P}\mathopen{}\mathclose{{}\left(\bigcup_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\bigcup_{d\in[D]}\mathopen{}\mathclose{{}\left\{\sum_{l=k_{b,d}^{\mathrm{v}}+1}^{k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil}\max_{b\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{b}+\text{1}_{B_{l}^{\mathrm{v}}}(b)\,\tau_{l}(\mathrm{v})\}-\sum_{l=k_{b,d}^{\mathrm{v}}+1}^{k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right\}}\right\}>\mathrm{C}\lceil T^{\alpha}\rceil^{\kappa}}\right)$
$\quad\leqslant\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\sum_{d\in[D]}\mathbb{P}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\{\sum_{l=k_{b,d}^{\mathrm{v}}+1}^{k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil}\max_{b\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{b}+\text{1}_{B_{l}^{\mathrm{v}}}(b)\,\tau_{l}(\mathrm{v})\}-\sum_{l=k_{b,d}^{\mathrm{v}}+1}^{k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right\}>\mathrm{C}\lceil T^{\alpha}\rceil^{\kappa}}\right\}}\right)$
$\quad\leqslant\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\sum_{d\in[D]}1/\lceil T^{\alpha}\rceil^{\zeta}$
$\quad\leqslant\mathrm{B}\,D/T^{\alpha\zeta}\;,$
where the first inequality holds with two union bounds, the second inequality holds because any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ satisfy 4.3 and the last inequality holds because $\text{Card}(\mathrm{C}(\mathrm{v}))=\mathrm{B}$. Considering the opposite event, we obtain that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, we have that the event
$\mathopen{}\mathclose{{}\left\{\bigcap_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\bigcap_{d\in[D]}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\lceil T^{\alpha}\rceil],\mathrm{C},\kappa)}\right\}$
holds with probability at least $1-\mathrm{B}D/T^{\alpha\zeta}$. Finally, the definition of $\mathcal{E}^{\mathrm{w}}_{\text{good}}$ gives the first line of the lemma. ∎
Lemma B.4.
Consider a player $\mathrm{v}\in\mathrm{V}$ who runs $D$ batches of binary search on action $b\in\mathcal{A}$ for all his children $\mathrm{C}(\mathrm{v})$. Assume that any child $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ satisfies 4.3 with parameters $(\mathrm{W}_{\mathrm{w}},\mathrm{C}_{\mathrm{w}},\kappa_{\mathrm{w}},\zeta_{\mathrm{w}})$ such that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, we have $\mathrm{W}_{\mathrm{w}}\leqslant\mathrm{W},\mathrm{C}_{\mathrm{w}}\leqslant\mathrm{W},\kappa_{\mathrm{w}}\leqslant\mathrm{W},\zeta_{\mathrm{w}}\geqslant\zeta$. Consider some action $b\in\mathcal{A}$ such that $D$ payment exploration batches of length $\tilde{T}$ are run on arm $b$ $D\in\{1,\ldots,\lceil\log_{2}T^{\beta}\rceil\}$ for any player $\mathrm{w}\in\mathrm{C}(\mathrm{v})$. Then
$\bigcap_{d\in[D]}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)\subseteq\{\tau^{\star}_{b}(\mathrm{w})\in[\underline{\tau}^{\mathrm{w}}_{b}(D),\bar{\tau}^{\mathrm{w}}_{b}(D)]\}\;,$
and the probability of these events is at least $1-\lceil\log T^{\beta}\rceil/T^{\alpha\zeta}$. We also have that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$
$\{|\bar{\tau}^{\mathrm{w}}_{b}(D)-\underline{\tau}^{\mathrm{w}}_{b}(D)|\leqslant 1/2^{D}+2/T^{\beta}\}\text{ holds almost surely}\;.$
Proof of Lemma B.4.
Suppose that the conditions of the lemma hold and consider a node $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ as well as some arm $b\in\mathcal{A}$.
The proof is done by induction on the number of binary search batches $D$ that are run on $b$ and player $\mathrm{w}$, such that $\bigcap_{d\in[D]}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)\subseteq\{\tau^{\star}_{b}(\mathrm{w})\in[\underline{\tau}^{\mathrm{w}}_{b}(D),\bar{\tau}^{\mathrm{w}}_{b}(D)]\}$.
The initialisation holds since $\underline{\tau}^{\mathrm{w}}_{a}(0)=0$, $\bar{\tau}^{\mathrm{w}}_{a}(0)=1$ and $\tau^{\star}_{b}(\mathrm{w})=\max_{b^{\prime}\in\mathcal{A}}\mu^{\star,\mathrm{w}}_{b^{\prime}}-\mu^{\star,\mathrm{w}}_{b}\in[0,1]$ with probability $1$ since $\max_{b^{\prime}\in\mathcal{A}}\mu^{\star,\mathrm{w}}_{b^{\prime}}\in[0,1]$ and $\mu^{\star,\mathrm{w}}_{b}\in[0,1]$.
Suppose that the property is true for some integer $D<\lceil\log T^{\beta}\rceil$ and that we have run one more binary search on arm $b$. We have
$\tau^{\text{mid},\mathrm{w}}_{b}(D+1)=\frac{\bar{\tau}^{\mathrm{w}}_{b}(D)+\underline{\tau}^{\mathrm{w}}_{b}(D)}{2}\;,$
[…]
Therefore, for any number $D$ of payment exploration batches run on arm $b$, we have that $\tau^{\star}_{b}(\mathrm{w})\in[\underline{\tau}^{\mathrm{w}}_{b}(D),\bar{\tau}^{\mathrm{w}}_{b}(D)]$ with probability at least $1-D/T^{\alpha\zeta}$. Lemma B.3 guarantees that $\bigcap_{d\in[D]}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)$ holds with probability at least $1-D/T^{\alpha\zeta}$.
[…]
$u_{D+1}^{\mathrm{w}}=\tau^{\text{mid},\mathrm{w}}_{b}(D+1)+1/T^{\beta}-\underline{\tau}^{\mathrm{w}}_{b}(D)=\frac{\bar{\tau}^{\mathrm{w}}_{b}(D)+\underline{\tau}^{\mathrm{w}}_{b}(D)}{2}-\underline{\tau}^{\mathrm{w}}_{b}(D)+1/T^{\beta}=u_{D}^{\mathrm{w}}/2+1/T^{\beta}\;.$
[…]
$u_{D+1}^{\mathrm{w}}=\bar{\tau}^{\mathrm{w}}_{b}(D)-(\tau^{\text{mid},\mathrm{w}}_{b}(D+1)-1/T^{\beta})=\bar{\tau}^{\mathrm{w}}_{b}(D)-\frac{\bar{\tau}^{\mathrm{w}}_{b}(D)+\underline{\tau}^{\mathrm{w}}_{b}(D)}{2}+1/T^{\beta}=u_{D}^{\mathrm{w}}/2+1/T^{\beta}\;.$
Thus, $(u_{D}^{\mathrm{w}})_{D\geqslant 1}$ is an arithmetico-geometric sequence defined by $u_{D+1}^{\mathrm{w}}=u_{D}^{\mathrm{w}}/2+1/T^{\beta}$ with an initial term $u_{1}^{\mathrm{w}}=1$. Writing $r=1/T^{\beta}/(1-1/2)=2/T^{\beta}$, we obtain that
$|\bar{\tau}^{\mathrm{w}}_{b}(D)-\underline{\tau}^{\mathrm{w}}_{b}(D)|=u_{D}^{\mathrm{w}}=1/2^{D}(1-r)+r=1/2^{D}(1-2/T^{\beta})+2/T^{\beta}\leqslant 1/2^{D}+2/T^{\beta}\;,$
for any $D\in\{1,\ldots,\lceil\log T^{\beta}\rceil\}$, hence the result. ∎
Proposition B.5.
Consider a player $\mathrm{v}\in\mathrm{V}$ such that any children $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ satisfies 4.3 with parameters $(\mathrm{W}_{\mathrm{w}},\mathrm{C}_{\mathrm{w}},\kappa_{\mathrm{w}},\zeta_{\mathrm{w}})$ and $\mathrm{W}_{\mathrm{w}}\leqslant\mathrm{W},\mathrm{C}_{\mathrm{w}}\leqslant\mathrm{C},\kappa_{\mathrm{w}}\leqslant\kappa,\zeta_{\mathrm{w}}\geqslant\zeta$. Suppose that 3.1 holds and that $\mathrm{v}$ runs Algorithm 2 with parameters $\mathrm{W},\mathrm{C},\kappa,\alpha,\beta,\eta$ on any action $b\in\mathrm{C}(\mathrm{v})$ for any player $\mathrm{w}\in\mathrm{C}(\mathrm{v})$. In that case, we have that the event
for $\text{for any }\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A},$
$\hat{\tau}_{b}(\mathrm{w})-4/T^{\beta}-\mathrm{C}\,\mathrm{B}\,T^{-\eta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\hat{\tau}_{b}(\mathrm{w})\;,$
holds with probability at least $1-K\mathrm{B}\lceil\log T^{\beta}\rceil/T^{\alpha\zeta}$.
Proof of Proposition B.5.
We start the proof by considering the opposite event and applying an union bound, following the definition $\hat{\tau}_{b}(\mathrm{w})=\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)+1/T^{\beta}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}$
(for $\mathbb{P}\mathopen{}\mathclose{{}\left(\text{for any }\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A},\hat{\tau}_{b}(\mathrm{w})-4/T^{\beta}-\mathrm{C}\,\mathrm{B}\,T^{-\eta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\hat{\tau}_{b}(\mathrm{w})}\right)$
$\quad\geqslant 1-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\sum_{b\in\mathcal{A}}\mathbb{P}\mathopen{}\mathclose{{}\left(\{\hat{\tau}_{b}(\mathrm{w})-4/T^{\beta}-\mathrm{C}\,\mathrm{B}\,T^{-\eta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\hat{\tau}_{b}(\mathrm{w})\}^{\mathrm{c}}}\right)$
$\quad=1-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\sum_{b\in\mathcal{A}}\mathbb{P}\mathopen{}\mathclose{{}\left(\{\bar{\tau}^{\mathrm{w}}_{b}-3/T^{\beta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\bar{\tau}^{\mathrm{w}}_{b}+1/T^{\beta}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}\}^{\mathrm{c}}}\right)\;,$
by definition of $\hat{\tau}_{b}(\mathrm{w})$ from (11). If $\bigcap_{d\in\lceil\log T^{\beta}\rceil}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)$ holds, then Lemma B.4 gives that $\tau^{\star}_{b}(\mathrm{w})\in[\underline{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil),\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)]$, and the second part of the lemma gives that
$|\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)-\underline{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)|\leqslant 1/2^{\lceil\log T^{\beta}\rceil}+2/T^{\beta}\leqslant 1/T^{\beta}+2/T^{\beta}=3/T^{\beta}\;.$
In that case, we can write
(for $\mathbb{P}\mathopen{}\mathclose{{}\left(\text{for any }\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A},\hat{\tau}_{b}(\mathrm{w})-4/T^{\beta}-\mathrm{C}\,\mathrm{B}\,T^{-\eta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\hat{\tau}_{b}(\mathrm{w})}\right)$
$\quad\geqslant 1-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\sum_{b\in\mathcal{A}}\mathbb{P}\mathopen{}\mathclose{{}\left(\{\underline{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\bar{\tau}^{\mathrm{w}}_{b}(\lceil\log T^{\beta}\rceil)\}^{\mathrm{c}}}\right)\;.$
However, Lemma B.3 ensures that $\bigcap_{d\in\lceil\log T^{\beta}\rceil}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,\ldots,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)$ happens with probability at least $1-\lceil\log T^{\beta}\rceil/T^{\alpha\zeta}$. Therefore,
(for $\mathbb{P}\mathopen{}\mathclose{{}\left(\text{for any }\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A},\hat{\tau}_{b}(\mathrm{w})-4/T^{\beta}-\mathrm{C}\,\mathrm{B}\,T^{-\eta}\leqslant\tau^{\star}_{b}(\mathrm{w})\leqslant\hat{\tau}_{b}(\mathrm{w})}\right)$
$\quad\geqslant 1-\mathrm{B}\,K\,\lceil\log T^{\beta}\rceil/T^{\alpha\zeta}\;.$
Proof of Lemma 4.2.
We first write our regret on the window $\{s+1,\ldots,s+t\}$ depending on whether all the children accepted the contract or not, following
$\mathfrak{R}^{\mathrm{v}}(s,t)$ $=\sum_{l=s}^{t}\max_{a,b^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})}\right\}-\sum_{l=s}^{t}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{B_{l}^{\mathrm{w}}}(A^{\mathrm{w}}_{l})\tau_{l}(\mathrm{w})}\right\}$
$=\sum_{l=s}^{t}\max_{a,b^{\mathrm{C}(\mathrm{v})}}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})}\right\}-\sum_{l=s}^{t}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w})}\right\}$
$\quad+\sum_{l=s}^{t}\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\text{1}_{B_{l}^{\mathrm{w}}}(A^{\mathrm{w}}_{l})\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w}))+\sum_{l=s}^{t}(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})}))$
$\leqslant\sum_{l=s}^{t}\mathopen{}\mathclose{{}\left\{\max_{a,b^{\mathrm{C}(\mathrm{v})}}\{\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}-(\mu^{\star,\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v}))}\right\}$
$\quad+\sum_{l=s}^{t}\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w}))_{+}+\sum_{l=s}^{t}(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})}))_{+}$
$\leqslant\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}(s,t)+\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}(s,t)+\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}(s,t)\;,$
hence the result. ∎
Lemma B.6.
Considering the setup of Lemma 4.2, we have that
$\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}(s,t)\leqslant\mathrm{B}\,(t-s+1)\max_{\underset{l\in\{s,\ldots,t\}}{\mathrm{w}\in\mathrm{C}(\mathrm{v})}}|\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w})|\;\;\text{ and }\;\;\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}(s,t)\leqslant\text{Card}\mathopen{}\mathclose{{}\left\{l\in\{s,\cdots,t\}\colon A_{l}^{\mathrm{C}(\mathrm{v})}\neq B_{l}^{\mathrm{C}(\mathrm{v})}}\right\}\;.$
Proof of Lemma B.6.
We start with the payment regret, which can be written
$\mathfrak{R}_{\mathrm{pay}}^{\mathrm{v}}(s,t)$ $=\sum_{l=s}^{t}\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w}))_{+}$
$\leqslant\sum_{l=s+1}^{s+t}\mathrm{B}\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}|\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w})|$
$\leqslant\sum_{l=s+1}^{s+t}\mathrm{B}\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\max_{l\in\{s+1,\ldots,s+t\}}|\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w})|$
$\leqslant\mathrm{B}\,t\max_{\underset{l\in\{s+1,\ldots,s+t\}}{\mathrm{w}\in\mathrm{C}(\mathrm{v})}}|\tau_{l}(\mathrm{w})-\tau^{\star}_{B_{l}^{\mathrm{w}}}(\mathrm{w})|\;.$
We can now decompose the deviation regret, following
$\mathfrak{R}_{\mathrm{dev}}^{\mathrm{v}}(s,t)$ $=\sum_{l=s}^{t}(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})}))_{+}$
$=\sum_{l=s}^{t}\text{1}_{B_{l}^{\mathrm{C}(\mathrm{v})}}(A_{l}^{\mathrm{C}(\mathrm{v})})\underbrace{(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})}))_{+}}_{\leqslant 1}$
$\leqslant\sum_{l=s}^{t}\text{1}_{B_{l}^{\mathrm{C}(\mathrm{v})}}(A_{l}^{\mathrm{C}(\mathrm{v})})$
$=\text{Card}\mathopen{}\mathclose{{}\left\{l\in\{s,\cdots,t\}\colon A_{l}^{\mathrm{C}(\mathrm{v})}\neq B_{l}^{\mathrm{C}(\mathrm{v})}}\right\}\;.$
Proof of Lemma 4.4.
Consider the setup of the lemma, as well as some set of steps $\{s+1,\ldots,s+t\}\subseteq\{\Psi_{K+1}^{\mathrm{v}}+1,\ldots,T\}$. We define the following ”nice” event
$\mathcal{E}=$ $\bigcap_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\bigcap_{d\in\lceil\log T^{\beta}\rceil}\bigcap_{b\in\mathcal{A}}\mathcal{E}^{\mathrm{w}}_{\text{good}}([k_{b,d}^{\mathrm{v}}+1,k_{b,d}^{\mathrm{v}}+\tilde{T}],\mathrm{C},\kappa)$
$\quad\bigcap_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathopen{}\mathclose{{}\left\{\sum_{l=s+1}^{s+t}\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{w}}_{a}+\text{1}_{B_{l}^{\mathrm{w}}}(a)\,\tau_{l}(\mathrm{w})\}-\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{w}}(A_{l}^{\mathrm{w}},B_{l}^{\mathrm{C}(\mathrm{w})})+\text{1}_{B_{l}^{\mathrm{w}}}(A_{l}^{\mathrm{w}})\,\tau_{l}(\mathrm{w})}\right\}\leqslant\mathrm{C}t^{\kappa}}\right\}\;,$
and a straight computation gives that
$\mathbb{P}(\mathcal{E})$ $\geqslant 1-\mathrm{B}\,K\,\lceil\log T^{\beta}\rceil/T^{\alpha\zeta}-\mathrm{B}t^{-\zeta}$
$\geqslant 1-\mathrm{B}\,K\,\lceil\log T^{\beta}\rceil/t^{\alpha\zeta}-\mathrm{B}/t^{-\alpha\zeta}$
$\geqslant 1-\max\mathopen{}\mathclose{{}\left\{\frac{2\,\mathrm{B}\,K\lceil\log T^{\beta}\rceil}{T^{\alpha\zeta}},\frac{2\,\mathrm{B}}{t^{\zeta}}}\right\}$
$\geqslant 1-\max\mathopen{}\mathclose{{}\left\{\frac{1}{T^{\alpha\zeta-\varepsilon}},\frac{2\,\mathrm{B}}{t^{\zeta}}}\right\}$
$\geqslant 1-1/t^{\alpha\zeta-\varepsilon}\;,$
for $\varepsilon$ is defined as
$\varepsilon=\log(4\mathrm{B}K\log T)/\log(T)\;,$ (22)
and $t\geqslant(2\mathrm{B})^{\zeta-\alpha\zeta+\varepsilon}$ with $\zeta-\alpha\zeta+\varepsilon>0$. We suppose that $\mathcal{E}$ holds for the rest of the proof. Based on the children $\mathrm{w}\in\mathrm{C}(\mathrm{v})$ time to learn, we can define player $\mathrm{v}$’s time to learn, following $\mathrm{W}_{\mathrm{v}}=\Psi_{K+1}^{\mathrm{v}}+\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{W}_{\mathrm{w}}$, and the following sets
$I_{\{s+1,\ldots,s+t\}}^{\mathrm{w}}$ =$=\{l\in\{s+1,\ldots,s+t\}\text{ such that }A_{l}^{\mathrm{w}}=B_{l}^{\mathrm{w}}\}\;,$
$J_{\{s+1,\ldots,s+t\}}^{\mathrm{w}}$ =$=\{l\in\{s+1,\ldots,s+t\}\text{ such that }A_{l}^{\mathrm{w}}\neq B_{l}^{\mathrm{w}}\}\;,$
as well as
$I^{\max}_{\{s+1,\ldots,s+t\}}(\mathrm{v})=\{l\in\{s+1,\ldots,s+t\}\text{ such that }l\in I_{\{s+1,\ldots,s+t\}}^{\mathrm{w}}\text{ for any }\mathrm{w}\in\mathrm{C}(\mathrm{v})\}\;,$
$J^{\max}(\mathrm{v})=\{l\in\{s+1,\ldots,s+t\}\text{ such that }l\in J_{\{s+1,\ldots,s+t\}}^{\mathrm{w}}\text{ for some }\mathrm{w}\in\mathrm{C}(\mathrm{v})\}\;.$
Note that $I_{\{s+1,\ldots,s+t\}}\cup J_{\{s+1,\ldots,s+t\}}=\{s+1,\ldots,s+t\}$ and $I^{\max}_{\{s+1,\ldots,s+t\}}\cup J^{\max}_{\{s+1,\ldots,s+t\}}=\{s+1,\ldots,s+t\}$. Since $\mathcal{E}$ holds, we have that for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, by definition of $\hat{\tau}_{b}(\mathrm{w})$
$\mu^{\star,\mathrm{w}}_{b}+\hat{\tau}_{b}(\mathrm{w})>\,\mu^{\star,\mathrm{w}}_{b}+\max_{b^{\prime\prime}\in\mathcal{A}}\mu^{\star,\mathrm{w}}_{b^{\prime\prime}}-\mu^{\star,\mathrm{w}}_{b}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}\,\geqslant\,\mu^{\star,\mathrm{w}}_{b^{\prime}}+\mathrm{C}\,\mathrm{B}\,T^{-\eta}\;,$
for any $b^{\prime}\in\mathcal{A}$ such that $b^{\prime}\neq b$. Therefore, we have that $B_{t}^{\mathrm{w}}=\operatorname*{\arg\!\max}_{b\in\mathcal{A}}\{\mu^{\star,\mathrm{w}}_{b}+\text{1}_{B_{t}^{\mathrm{w}}}(b)\,\hat{\tau}_{b}(\mathrm{w})\}$, as well as
$(\mu^{\star,\mathrm{w}}_{B_{t}^{\mathrm{w}}}+\text{1}_{B_{t}^{\mathrm{w}}}(B_{t}^{\mathrm{w}})\,\hat{\tau}_{B_{t}^{\mathrm{w}}})-\max_{\underset{b\neq B_{t}^{\mathrm{w}}}{b\in\mathcal{A}}}\{\mu^{\star,\mathrm{w}}_{b}+\text{1}_{B_{t}^{\mathrm{w}}}(b)\,\hat{\tau}_{B_{t}^{\mathrm{w}}}\}\geqslant\mathrm{C}\,\mathrm{B}\,T^{-\eta}\;.$ (23)
Thus, for any step $l\in\{s+1,\ldots,s+t\}$, (23) implies that the reward gap between $B_{t}^{\mathrm{w}}$ and any other action $b\in\mathcal{A}$ is at least $\mathrm{C}\,\mathrm{B}\,T^{-\eta}$, and hence
$\text{Card}(J_{\{s+1,\ldots,s+t\}}^{\mathrm{w}})\,\mathrm{C}\,T^{-\eta}\,\mathrm{B}$ $\leqslant\sum_{l\in\{s+1,\ldots,s+t\}}\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{a}+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}-\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right\}$
$\leqslant\mathrm{C}\,t^{\kappa}\;,$
by 4.3. We can conclude that $\text{Card}(J_{\{s+1,\ldots,s+t\}}^{\mathrm{w}})\leqslant t^{\kappa}/(T^{-\eta}\,\cdot\mathrm{B})\leqslant t^{\kappa+\eta}/\mathrm{B}$ for any $\mathrm{w}\in\mathrm{C}(\mathrm{v})$, which gives that
$\text{Card}(J^{\max}_{\{s+1,\ldots,s+t\}})\leqslant\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{Card}(J_{\{s+1,\ldots,s+t\}}^{\mathrm{w}})\leqslant t^{\kappa+\eta}\;.$ (24)
We can now decompose the quantity that we want to bound
$\sum_{l=s+1}^{s+t}\mathopen{}\mathclose{{}\left\{\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{a}+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}-\mathopen{}\mathclose{{}\left(\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right)}\right\}\leqslant 2\,\text{Card}(J^{\max}_{\{s+1,\ldots,s+t\}})$
$\quad+\sum_{l\in I^{\max}_{\in\{s+1,\ldots,s+t\}}}\mathopen{}\mathclose{{}\left\{\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{a}+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}-\mathopen{}\mathclose{{}\left(\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right)}\right\}$
$\leqslant 2\,t^{\kappa+\eta}+\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\max_{a,b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}\times\mathcal{A}^{\mathrm{B}}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\tau_{l}(\mathrm{v})}\right\}$
$\quad-\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{A_{l}^{\mathrm{w}}}(\mathrm{w})}\right)$
$=2\,t^{\kappa+\eta}+\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\max_{a,b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}\times\mathcal{A}^{\mathrm{B}}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\tau^{\star}_{b^{\mathrm{w}}}(\mathrm{w})+\mathrm{C}\,\mathrm{B}\,T^{-\eta}+4T^{-\beta})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\tau_{l}(\mathrm{v})}\right\}$
$\quad-\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}(\tau^{\star}_{A_{l}^{\mathrm{w}}}(\mathrm{w})+\mathrm{C}\,\mathrm{B}\,T^{-\eta})}\right)+\,4\,\mathrm{B}\,t\,T^{-\beta}$
$\leqslant 2\,t^{\kappa+\eta}+\,4\,\mathrm{B}\,t^{1-\beta}+\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\max_{a,b^{\mathrm{C}(\mathrm{v})}\in\mathcal{A}\times\mathcal{A}^{\mathrm{B}}}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{b^{\mathrm{w}}}(\mathrm{w})+\text{1}_{B_{l}^{\mathrm{v}}}(a)\tau_{l}(\mathrm{v})}\right\}$
$\quad-\sum_{l\in I^{\max}_{\{s+1,\ldots,s+t\}}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{A_{l}^{\mathrm{w}}}(\mathrm{w})}\right)\;,$
thanks to Proposition B.5. But player $\mathrm{v}$ plays on a stochastic bandit with reward and observed feedback $Z_{l}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{C}(\mathrm{v})})=X_{l}(A_{l}^{\mathrm{v}},A_{l}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{A_{t}^{\mathrm{w}}}(\mathrm{w})+\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{C}(\mathrm{v})}$ as long as $l\in I^{\max}_{\{s+1,\ldots,s+t\}}$. Hence, Scheid et al. (2024a, Proposition 2) gives us that
$\sum_{l=s+1}^{s+t}\mathopen{}\mathclose{{}\left\{\max_{a\in\mathcal{A}}\{\mu^{\star,\mathrm{v}}_{a}+\text{1}_{B_{l}^{\mathrm{v}}}(a)\,\tau_{l}(\mathrm{v})\}-\mathopen{}\mathclose{{}\left(\mu^{\mathrm{v}}(A_{l}^{\mathrm{v}},B_{l}^{\mathrm{C}(\mathrm{v})})-\text{1}_{B_{l}^{\mathrm{v}}}(A_{l}^{\mathrm{v}})\,\tau_{l}(\mathrm{v})}\right)}\right\}$ $\leqslant 2(t^{\kappa+\eta}+t^{1-\beta})$
$\quad+8\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3})}t^{1/2},$
since player $\mathrm{v}$ plays on a $K^{\mathrm{B}+1}$ bandit instance. Hence, $\mathrm{v}$ satisfies 4.3 with parameters $(\mathrm{W}+K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,10\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3})},\max\{\kappa+\eta,1-\beta,1/2\},\alpha\zeta-\varepsilon)$, where $\varepsilon=\log(4\mathrm{B}K\log T)/\log(T)$. ∎
Proof of Lemma 4.5.
We do the proof by induction, starting with the layer $d=2$: the layer of players interacting with the leafs. Scheid et al. (2024a, Proposition 2) gives us that Algorithm 3 for the leafs without offering any transfer satisfies 4.3 with parameters $(0,8\sqrt{K\log(KT^{3})},1/2,2)$. Thus, following (14), we choose parameters $\eta_{2}=1/4,\alpha_{2}=(D+1)/2D,\beta_{2}=1/4$ for any $\mathrm{v}$ in layer $d=2$. $\kappa_{1}+\eta_{2}=3/4,1-\beta_{2}=3/4$. Therefore, Lemma 4.4 ensures that $\mathrm{v}$ satisfies 4.3 with parameters $(K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,10\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3}}),3/4,4/3-\varepsilon)=((d-1)\,K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,10\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3}}),1-1/2d,\zeta_{d})$ for $d=2$. Thus, our initialization holds.
Suppose that the property is true for some node $\mathrm{w}$ located at depth $d$ and consider a node $\mathrm{v}=\mathrm{P}(\mathrm{w})$ at depth $d+1$. $\mathrm{v}$ runs COBRA with parameters from (14). It is important to check if (13) is satisfied. Note that for any $d\in\{1,\ldots,D\}$, we have that
$\frac{\beta_{d+1}}{\alpha_{d+1}}=\frac{1}{2(d+1)}\times\frac{D(d+1)}{(D+1)d}=\frac{D}{D+1}\times\frac{1}{2d}\;,$
while we have that
$1-\kappa_{d}=1-1+\frac{1}{2d}=\frac{1}{2d}>\beta_{d+1}/\alpha_{d+1}\;,$
hence $\alpha_{d+1},\beta_{d+1}$ are acceptable parameters for a player at depth $d$ since they satisfy $\beta_{d+1}/\alpha_{d+1}<1-\kappa_{d}$. We now have that
$\kappa_{d+1}$ $=\max(1/2,\kappa_{d}+\eta_{d+1},1-\beta_{d+1})$
$=\max\mathopen{}\mathclose{{}\left\{1/2,1-1/(2\,\cdot d)+1/\{2\,d\,(d+1)\},1-1/2(d+1)}\right\}$
$=\max\mathopen{}\mathclose{{}\left\{1-\frac{1}{2d}\mathopen{}\mathclose{{}\left(1-\frac{1}{d+1}}\right),1-\frac{1}{2(d+1)}}\right\}\;\;\;\text{for any d}\geqslant 3$
$=\max\mathopen{}\mathclose{{}\left\{1-\frac{1}{2d}\frac{d+1-1}{d+1},1-\frac{1}{2(d+1)}}\right\}$
$=1-\frac{1}{2(d+1)}\;.$
Now observe that following Lemma 4.5 with $\varepsilon$ defined in (22) as
$\varepsilon=\log(4\mathrm{B}K\log T)/\log(T)\;,$
we have that $\zeta_{d+1}=\alpha_{d+1}\zeta_{d}-\varepsilon$ and Lemma B.7 ensures that $\zeta_{d}\geqslant 2/d-d\varepsilon$. Since we assumed that $\log T\geqslant D^{2}\log(4\mathrm{B}K\log T)$, we have for any $d\in\{1,\ldots,D\},\log T\geqslant d^{2}\log(4\mathrm{B}K\log T)$, which implies $1/d\geqslant d\log(4\mathrm{B}K\log T)/\log T=d\varepsilon$, and hence we obtain that $\zeta_{d}\geqslant 1/d$. ∎
Lemma B.7.
If $(\zeta_{d})_{d\geqslant 1}$ is defined as $\zeta_{1}=2$ and for any $d\geqslant 1$, $\zeta_{d+1}=\alpha_{d+1}\zeta_{d}-\varepsilon$, for $\varepsilon\in(0,1)$, then we have that
$\zeta_{d}\geqslant\frac{2}{d}-\varepsilon\mathopen{}\mathclose{{}\left(\frac{d+1}{2}-\frac{1}{d}}\right)\geqslant\frac{2}{d}-d\,\varepsilon\;.$
Proof of Lemma B.7.
First define the sequence $(v_{d})_{d\geqslant 1}$ as $v_{1}=2$ and then $v_{d+1}=dv_{d}/(d+1)-\varepsilon$. We show by induction that for any $d\geqslant 1,\zeta_{d}\geqslant v_{d}$. The definition of $\zeta_{1}=2=v_{d}$ and now assuming that $\zeta_{d}\geqslant v_{d}$ for some $d\geqslant 1$, we obtain
$\zeta_{d+1}$ $=\frac{(D+1)(d-1)}{D\times d}\zeta_{d}-\varepsilon$
$\geqslant\frac{d}{d+1}\zeta_{d}-\varepsilon$
$\geqslant\frac{d}{d+1}v_{d}-\varepsilon\;\;\;\text{by our induction assumption}$
$=v_{d+1}\;,$
by definition. Some algebra gives that $v_{d}=\frac{2}{d}+\varepsilon\mathopen{}\mathclose{{}\left(\frac{1}{d}-\frac{d+1}{2}}\right)\geqslant\frac{2}{d}-d\,\varepsilon$, hence the result. ∎
Proof of Theorem 4.6.
Consider player $\mathrm{v}\in\mathrm{V}$ located at depth $d$. Since all the nodes run COBRA, Lemma 4.4 gives us that $\mathrm{C}(\mathrm{v})$ satisfy 4.3 with parameters $(\mathrm{W}_{\mathrm{w}},\mathrm{C}_{\mathrm{w}},\kappa_{\mathrm{w}},\zeta_{\mathrm{w}})$ and we have that $\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\mathrm{W}_{\mathrm{w}}\leqslant K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil$, $\mathrm{C}_{\mathrm{w}}\leqslant 10\sqrt{K^{\mathrm{B}+1}\log(K^{\mathrm{B}+1}T^{3})}$. Also, $\kappa_{d}=1-1/(2d-2)$ and $\zeta_{d}\geqslant 1/d$. Player $\mathrm{v}$ ’s instantaneous regret is always bounded by $2$ almost surely, hence we can write
$\mathfrak{R}^{\mathrm{v}}(T)$ $\leqslant 2\,d\,K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil+\sum_{s=\mathrm{W}+K\lceil T^{\alpha}\rceil\lceil\log(T^{\beta})}^{T}\max_{a,(b^{\mathrm{C}(\mathrm{v})})}\mathopen{}\mathclose{{}\left\{\mu^{\mathrm{v}}(a,b^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(a)\,\tau_{t}(\mathrm{v})}\right\}$
$\quad-\sum_{s=\mathrm{W}+K\lceil T^{\alpha}\rceil\lceil\log(T^{\beta})}^{T}\mathopen{}\mathclose{{}\left\{\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\,\tau_{t}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{B_{t}^{\mathrm{w}}}(A^{\mathrm{w}}_{t})\tau_{t}(\mathrm{w})}\right\}$
Using Lemma 4.2, we can decompose $\mathfrak{R}^{\mathrm{v}}$ as follows
$\mathfrak{R}^{\mathrm{v}}(T)$ $\leqslant 2\,d\,K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil+\mathfrak{R}_{\mathrm{ac}}^{\mathrm{v}}((d+1)K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,T)$
$\quad+2\,\text{Card}(s\in\{1+dK\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,\ldots,T\}\colon A_{s}^{\mathrm{C}(\mathrm{v})}\neq B_{s}^{\mathrm{C}(\mathrm{v})})$
$\quad+\mathrm{B}\,(T-dK\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil)\max_{\underset{s\geqslant(d+1)K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil}{\mathrm{w}\in\mathrm{C}(\mathrm{v})}}|\hat{\tau}_{B_{s}^{\mathrm{w}}}(\mathrm{w})-\tau^{\star}_{B_{s}^{\mathrm{w}}}(\mathrm{w})|$
$\leqslant 2\,d\,K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil+\mathrm{C}_{\mathrm{v}}(T-K\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil)^{\kappa_{\mathrm{v}}}+\mathrm{B}\,T\,\max_{b\in\mathcal{A},\mathrm{w}\in\mathrm{V}}|\hat{\tau}_{b}(\mathrm{w})-\tau^{\star}_{b}(\mathrm{w})|$
$\quad+2\,\text{Card}(s\in\{1+dK\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil,\ldots,T\}\colon A_{s}^{\mathrm{C}(\mathrm{v})}\neq B_{s}^{\mathrm{C}(\mathrm{v})})\;,$
with probability at least $1-1/T^{-\zeta_{d}}$ by Lemma 4.4. Following what is shown in (24), we have that for any $s\geqslant\mathrm{W}$
$\text{Card}(J^{\max}_{\{s+1,\ldots,s+t\}})\leqslant t^{\kappa_{d-1}+\eta_{d}}\;,$
and Proposition B.5 gives that
$\max_{\mathrm{w}\in\mathrm{C}(\mathrm{v}),b\in\mathcal{A}}|\hat{\tau}_{b}(\mathrm{w})-\tau^{\star}_{b}(\mathrm{w})|\leqslant 4/T^{\beta_{d}}+\mathrm{C}\,\mathrm{B}\,T^{-\eta_{d}}\;.$
Therefore, we can write, with probability at least $1-1/T^{\zeta_{d}}$
$\mathfrak{R}^{\mathrm{v}}(T)$ $\leqslant 2\,d\,K\lceil T^{\alpha_{d}}\rceil\lceil\log T^{\beta_{d}}\rceil+\mathrm{C}\,T^{\kappa_{d}}+2(T-d\,K\lceil T^{\alpha_{d}}\rceil\lceil\log T^{\beta_{d}}\rceil)^{\kappa_{d-1}+\eta_{d}}+\mathrm{B}\,T\,(4\,T^{-\beta_{d}}+\mathrm{C}\,\mathrm{B}\,T^{-\eta_{d}})$
$=\mathcal{O}(T^{\alpha_{d}}+T^{\kappa_{d-1}+\eta_{d}}+T^{1-\beta_{d}}+T^{1-\eta_{d}})$
$=\mathcal{O}(T^{1-\frac{1}{2d^{2}}})\;,$
by definition of the parameters in (14) and Lemma 4.4. For the comment that follows Theorem 4.6, we can write (since $\zeta_{d}\geqslant 1/d$)
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{v}}(T)}\right]\leqslant(1-1/T^{\zeta_{d}})\mathcal{O}(T^{1-\frac{1}{2d^{2}}})+2\times T/T^{\zeta_{d}}\leqslant\mathcal{O}(T^{1-\frac{1}{2d^{2}}})+\mathcal{O}(T^{1-\frac{1}{d}})=\mathcal{O}(T^{1-\frac{1}{2d^{2}}})\;,$
hence the fact that $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{v}}(T)}\right]=o(T)$. ∎
Proof of Corollary 4.7.
Following Lemma 4.1, we can write
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right\}=\sum_{t=1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})-\sum_{t=1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})\;.$
Since all the players run COBRA, writing $\mathrm{W}=(D+1)\,K\,\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil\geqslant\sum_{d=1}^{D}\max_{\mathrm{w}\in\mathrm{V}_{d}}\mathrm{W}_{\mathrm{w}}$, we can write
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right\}$
$\quad\leqslant 2\,\mathrm{W}+\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})\quad-\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}(\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v})-\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v}))\;,$
and reorganizing the terms $\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v})$ among the players allows us to write
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right\}$
$\quad\leqslant 2\,\mathrm{W}+\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})-\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\text{1}_{B_{t}^{\mathrm{w}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{w})}\right)$
$\leqslant 2\,\mathrm{W}+\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})-\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\hat{\tau}_{B_{t}^{\mathrm{w}}}(\mathrm{w})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{w})}\right)$
$\leqslant 2\mathrm{W}+\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})-\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left(\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})-\sum_{\mathrm{w}\in\mathrm{C}(\mathrm{v})}\tau^{\star}_{B_{t}^{\mathrm{w}}}(\mathrm{w})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v})}\right)$
$\quad\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left\{T\,\max_{b\in\mathcal{A},\mathrm{w}\in\mathrm{C}(\mathrm{v})}|\tau^{\star}_{b}(\mathrm{w})-\hat{\tau}_{b}(\mathrm{w})|+\text{Card}\mathopen{}\mathclose{{}\left(t\geqslant\mathrm{W}+1\colon A_{t}^{\mathrm{C}(\mathrm{v})}\neq B_{t}^{\mathrm{C}(\mathrm{v})}}\right)}\right\}$
$\leqslant 2\mathrm{W}+\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\max_{a^{\mathrm{v}}\in\mathcal{A}}\mu^{\star,\mathrm{v}}(a^{\mathrm{v}})-\sum_{t=\mathrm{W}+1}^{T}\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left(\mu^{\mathrm{v}}(A_{t}^{\mathrm{v}},B_{t}^{\mathrm{C}(\mathrm{v})})+\text{1}_{B_{t}^{\mathrm{v}}}(A_{t}^{\mathrm{v}})\tau_{t}(\mathrm{v})}\right)$
$\quad+\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left\{T\,(4/T^{\min_{d}\beta_{d}}+\mathrm{B}\,\max_{d}\mathrm{C}_{d}\,T^{-\min_{d}\eta_{d}})+T^{\max_{d}\kappa_{d}+\eta_{d}}}\right\}$
$\leqslant 2\,(D+1)\,K\,\lceil T^{\alpha}\rceil\lceil\log T^{\beta}\rceil+\sum_{\mathrm{v}\in\mathrm{V}}\mathopen{}\mathclose{{}\left(\mathfrak{R}^{\mathrm{v}}(T)+4\,T^{1-\min_{d}\beta_{d}}+\mathrm{B}\max_{d}\mathrm{C}_{d}\,T^{1-\min_{d}\eta_{d}}+T^{max_{d}\kappa_{d}+\eta_{d}}}\right)\;,$
with probability at least $1-\sum_{\mathrm{v}\in\mathrm{V}}1/T^{\zeta_{\mathrm{v}}}\to 0$ as $T\to+\infty$. Therefore, dividing by $T$, taking the expectation as in the proof of Theorem 4.6, and using Theorem 4.6 for the convergence of any $\mathfrak{R}^{\mathrm{v}}(T)$, we obtain that
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right\}/T\to 0\;,$
and therefore
$\sum_{t=1}^{T}\max_{(a^{\mathrm{v}})\in\mathcal{A}^{|\mathrm{V}|}}\mathopen{}\mathclose{{}\left\{\sum_{\mathrm{v}\in\mathrm{V}}\theta^{\mathrm{v}}(a^{\mathrm{v}},a^{\mathrm{C}(\mathrm{v})})-\theta^{\mathrm{v}}(A_{t}^{\mathrm{v}},A_{t}^{\mathrm{C}(\mathrm{v}})}\right\}=o(T)\;,$
hence our result which gives the convergence of the players’ global reward towards the optimum. ∎